### v2

# FormulaQA: A Question Answering Dataset for Formula-Based Numerical Reasoning

Abstract
The application of formulas is a fundamental ability of humans when addressing numerical reasoning problems. However, existing numerical reasoning datasets seldom explicitly indicate the formulas employed during the reasoning steps. To bridge this gap, we propose a question answering dataset for formula-based numerical reasoning called FormulaQA, from junior high school physics examinations.  We further conduct evaluations on LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thoughts methods and we explored the approach of using retrieval-augmented LLMs when providing an external formula database. We also fine-tune on smaller models with size not exceeding 2B. Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaQA.
FormulaQA: A Question Answering Dataset for
Formula-Based Numerical Reasoning
Xiao Li and Sichen Liu and Bolin Zhu and Yin Zhu and Yiwei Liu and Gong Cheng State Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China {xiaoli.nju,sichenliu,bolinzhu,yinzhu,ywliu}@smail.nju.edu.cn gcheng@nju.edu.cn
1 Introduction
Dataset Math23K-F MAWPS-F FormulaQA
# questions 23,162 2,373 5,420
# formulas (and variants) 51 (131) 18 (46) 271 (824)
# questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) 5,420 (100%)
Avg. # reasoning steps 1.16 1.01 2.37
Table 1: Statistics of Math23-F, MAWPS-F and our FormulaQA.
Figure 1: An example taken from FormulaQA. All the numerical values (including units) are highlighted in red font. In the formulas, we marked the calculated intermediate results with purple font. All formulas are highlighted in blue.
Numerical reasoning constitutes one of the significant forms within natural language reasoning Frieder et al. (2023). The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs OpenAI (2023); Touvron et al. (2023); Li et al. (2023c) and specialized datasets Wang et al. (2017); Dua et al. (2019); Amini et al. (2019); Cobbe et al. (2021a). Current datasets for numerical reasoning typically include simple, commonsense numerical questions that do not reflect the complexity of real-world problems. These datasets do not fully address the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process. This issue becomes particularly evident when LLMs meet hallucination Frieder et al. (2023); Bang et al. (2023). Consequently, one might naturally wonder that “What knowledge could I use to guide numerical reasoning process?”. Formulas just represent the type of knowledge that is often overlooked in research but frequently utilized in real-life applications.
Take a question from the GSM8K Cobbe et al. (2021a) dataset as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?”. This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula. However, in FormulaQA, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object. We annotated formulas for each question in FormulaQA. An example of FormulaQA is shown in Figure 1. The formula-based feature makes FormulaQA a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems. In fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge. However, existing datasets scarcely consider the incorporation of formulas into numerical reasoning.
Recently, Liu et al., 2023 constructed two formula-based datasets, Math23K-F and MAWPS-F. However, the formulas in these datasets primarily consist of commonsense formulas (such as total_amount = unit_amount $\times$ total_number), and only 33.5% and 38.4% of the questions in each dataset, respectively, require the use of formulas.
To address this gap, we constructed a question answering dataset for numerical reasoning that requires the use of formulas called FormulaQA. We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations. With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database. The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems. In Table 1, we compare FormulaQA with two existing formula-based datasets. In comparison to Math23K-F and MAWPS-F, FormulaQA contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas. Additionally, all questions in FormulaQA require the use of formulas. The higher average number of reasoning steps (2.37 vs. 1.16/1.01) implies that FormulaQA is more challenging and better suited for evaluating existing models as a multi-step reasoning task.
We evaluated LLMs ranging from 7B to >100B parameters, as well as fine-tuned smaller models such as mT5-base/large Xue et al. (2021) and Qwen-1.8B Bai et al. (2023) on FormulaQA. We also attempted to train an encoder for formula retrieval and experimented with retrieval-augmented generation models. Our evaluation findings show that the best existing models only achieve an accuracy of around 70%, which is not yet satisfactory, indicating that there is still significant room for exploration in formula-based numerical reasoning.
Our contributions can be summarized as follows:
  • We construct a formula-based numerical reasoning dataset FormulaQA, with comprehensive annotations for each question.
  • We conduct evaluations on LLMs of various sizes, fine-tuned smaller models, and retrieval-augmented generative models, establishing a strong baseline for future research.
Dataset and code are available on https://github.com/nju-websoft/FormulaQA under the Apache License 2.0.
2 Related Work
2.1 Numerical Reasoning Datasets
Numerical reasoning is one of the fundamental capabilities of natural language reasoning. The study of numerical reasoning in natural language has existed for several years. Numerous datasets, such as DROP Dua et al. (2019), GSM8K Cobbe et al. (2021b), TSQA Li et al. (2021) and MATH Hendrycks et al. (2021), have introduced natural language numerical reasoning. Another line of research focusing on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., the question) and require the generation of an arithmetic expression that can compute the answer. Representative datasets include MAWPS Koncel-Kedziorski et al. (2016), Math23K Wang et al. (2017), MathQA Amini et al. (2019), etc.
The recently introduced datasets Liu et al. (2023) Math23K-F and MAWPS-F require formulas for only 33.5% and 38.4% of the questions, respectively, and the formulas within these datasets are all simple commonsense formulas (e.g., total_cost=unit_cost $\times$ total_number). Our FormulaQA dataset collects questions from junior high school physics examinations, with each question accompanied by domain-specific formulas. In addition, we have also annotated a formula database for FormulaQA that can serve as an external knowledge base, used to assess retrieval-augmented systems.
2.2 Numerical Reasoning Methods
The methods for solving numerical reasoning have evolved from statistical approaches Hosseini et al. (2014); Kushman et al. (2014) to those based on rules and templates Shi et al. (2015); Wang et al. (2019) and further to methods based on deep learning models Gupta et al. (2019); Chen et al. (2022); Kim et al. (2022); Li et al. (2023a). In the past two years, with the rapid development of LLMs, LLMs have demonstrated strong capabilities in resolving numerical reasoning questions. Consequently, several methods aimed at enhancing the reasoning abilities of LLMs have been proposed, including the notable Chain of Thoughts (CoTs) method Wei et al. (2022), along with many subsequent variant approaches Kojima et al. (2022); Wang et al. (2022); Zhou et al. (2022); Li et al. (2023b).
We have established representative existing methods as baselines for FormulaQA, including fine-tuning on smaller-scale models with fewer than 2 billion parameters, such as mT5-base/large, and Qwen-1.8B. Additionally, we have applied zero-shot and few-shot CoTs prompting methods to LLMs ranging from 7 billion to over 100 billion parameters. Given that existing numerical reasoning methods scarcely consider retrieval-based approaches, as the existing numerical reasoning datasets mostly rely on implicit commonsense knowledge, in FormulaQA, we have trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning.
3 Dataset Construction
We collected raw questions from Chinese junior high school physics examinations from 2015 to the present, filtering out those that do not require formulas for solving. We then annotated the reasoning steps and corresponding formulas for each question. This process involved a combination of manual annotation and the assistance of LLMs to improve the efficiency of annotation. Each question included an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all parameters within the formulas. Finally, we compiled all the formulas we obtained and merged those expressing the same meaning to create a formula database. We describe this process to construct FormulaQA in detail below.
3.1 Preprocessing
We crawled 18,433 junior high school physics examination questions in China from 2015 to the present from public sources, including only those with free-text answers and excluding multiple-choice and true/false questions. Each original question contains a question text and an explanation text that includes the reasoning steps. We eliminated questions requiring diagrams.
Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical. Utilizing a rule-based approach, we extracted the final numeric answer, including its corresponding unit, from the explanation. It was observed that for 487 instances, the final answer had not been accurately extracted from the explanations, prompting us to perform manual computations and annotations to determine the correct final answers for these questions. Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
3.2 Formula Normalization
Original explanation.
[…]
Explanation with normalized formulas.
Calculating the degree of temperature increase in water: [Degree of water temperature increase] = [Final temperature] - [Initial temperature] = 60 ℃ - 20 ℃ = 40 ℃. The degree of water temperature increase = 40 ℃.
Calculating the heat absorbed by water: [Heat absorbed by water] = [Mass of water] $\times$ [Specific heat capacity of water] $\times$ [Degree of water temperature increase] = 50 kg $\times$ 4.2 $\times$ $10^{3}$ J/(kg·℃) $\times$ 40 ℃ = 8400000 J. The heat absorbed by water = 8400000 J.
The thermal efficiency of the water heater can be obtained from: [Thermal efficiency of the water heater] = [Heat absorbed by water] / [Total electrical energy consumed] $\times$ 100% = 8400000 J / ($1\times 10^{7}$ J) * 100% = 84%. The thermal efficiency of the water heater = 84%.
Answer = 84%
Table 2: Original explanation and explanation with normalized formulas (highlighted in blue).
We found that the reasoning steps (i.e. the explanation) in the obtained raw dataset lacked a normalized format and were expressed quite casually. Some formulas mixed parameter names (e.g., “mass of water”) and symbols (e.g., “$m_{water}$”), while others simply provided calculations in numerical form without parameter names or symbols. In order to ensure that explanations were based on a normalized form of formulas, we then normalized the formula annotations in the explanations. An example can be found in Table 2. In this process, we need to identify the formulas used within the original explanations and to correct any formatting issues. Manually undertaking such tasks required significant effort. However, since the process is not open-ended, but rather structured, methodical and verifiable, we ensured the accuracy and consistency of the numerical reasoning by applying the extraction of formulas from the explanations. Once the formulas were identified, we could calculate each step to ensure that the results aligned with the corresponding answers. Any redundant formulas or missing parameters could be detected, so we could introduce the assistance of LLMs here.
To enhance the efficiency of the annotation process, during the LLM annotation phase, we adopted a coarse-to-fine annotation approach with the help of a LLM [^1] During dataset constructions, we accessed Qwen-max via API (https://help.aliyun.com/zh/dashscope/developer-reference/quick-start). Qwen-max is a LLM with over 100 billion parameters and a strong capability in understanding Chinese.. We first prompted the LLM in a few-shot manner to generate accurate explanations of the reasoning process. Then, we used few-shot prompts to guide the LLM in correcting minor errors within the normalized explanations, including formatting errors in formula annotations and inaccuracies in the parameters used during computations. Both prompts can be found in Appendix A.1.1. Next, we will provide a detailed description of this process.
Initially, we introduced the question along with its original explanation and the corresponding answer to guide the LLM through few-shot prompting to revise the original explanation. We observed that the ability of the LLM to revise explanations towards normalized explanations remains satisfactory. To assess the correctness of the revised explanations, we extracted formulas from these explanations and then computed the answer using the numbat tool [^2] https://numbat.dev. Numbat is designed for scientific computations with support for physical dimensions and units.. In addition to providing explanations, we also required the LLM to present the values, symbols, and units of each parameter in the formulas in the form of a table. An example is shown in Figure 1.
At this stage, we used a program to check the correctness of the formula format in the explanations, including whether there were omissions in parameter names, parameter symbols, and corresponding units and these issues are all correctable. Therefore, if our program detected that the LLM had not successfully generated an accurate normalized explanation, we used few-shot prompting to identify and correct these specific errors. More details can be found in Appendix A.1.1. We observed that the questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, unclear question formulation, and so on. Some examples of these questions can be found in Appendix A.1.2. These questions have been removed from our dataset. Following this step, our dataset contains a remaining total of 5,420 questions.
3.3 Formula Database Construction
Our next step was to construct a unified formula repository for the entire dataset. Given that parameters in the same formula can be expressed differently across various problem contexts, for instance, the two formulas “[weight of water] = [mass of water] * [gravitational acceleration]” and “[weight] = [mass] * [gravitational acceleration]” both calculate the weight of an object, we need to merge these formulas into a single representation.
We divided the construction process of the formulas database into three steps: 1) Merge the formulas through symbolic rules. 2) Merge the formulas through semantics based method. 3) Manual review and error correction. In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
Step # Formulas
Before merging 12,906
After symbolic rules based merging 1,163
After semantics based merging 439
After manual review and error correction 272
Table 3: Changes in the number of formulas after each merging step.
Symbolic rules based merging.
In this step, we merged formulas through symbolic rules. Specifically, this was achieved by comparing the structure of the formulas and the symbols contained within them. Taking this as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}$=($b_{1}$+$c_{1}$)/$d_{1}$”, “$f_{2}:a_{2}$=($b_{2}$+$c_{2}$)/$d_{2}$” and “$f_{3}:b_{1}$=$a_{1}$*$d_{1}$-$c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by moving $b_{1}$ to the left side of the equation. Moreover, in physics, certain physical quantities are conventionally represented by specific symbols. For example, the mass of an object is often denoted by “$m$” and the density of an object is frequently represented by the symbol “$\rho$”. Subscripts are then used to distinguish which specific object the physical quantity refers to, such as “$\rho_{water}$” for the density of water. For any two formulas, we first computed all transformations of a formula to obtain a set of all variants of the formula. Then, we compared the formula structures in the two sets to determine if two formulas were structurally equivalent. If they shared the same structure, we then compared whether the symbols, with subscripts removed, were identical. If they were, we considered these two formulas to be mergeable. When merging, we retained the parameter with the shorter length from the two.
Semantics based merging.
In the symbolic rules based merging process, the semantic information of the parameter names was neglected. This led us to perform merges grounded on the semantics of the parameter names. For instance, two formulas that were not merged during the symbol fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged. We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, "density" and "density of water" are semantically similar). For formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^3] We used jieba: https://github.com/fxsjy/jieba.. When the two sets overlapped, the parameters were considered to have semantic connection, and the formulas became candidates for merging. Utilizing this approach, we identified a list of potential formula mergers and then consulted the LLM for a thorough evaluation of each pair. The prompts can be found in Appendix A.1.3.
Manual review and error correction.
Upon completing the aforementioned merging process, we manually inspected the correctness of the merge, rectified instances where errors occurred in the merging, and manually merged formulas that were overlooked by the LLM.
4 Experiments Setup
In this section, we explore several methods to address the questions within FormulaQA. These include prompting LLMs using zero-shot and few-shot chains of thoughts (CoTs, Wei et al., 2022; Kojima et al., 2022), fine-tuning smaller models with less than 2 billion parameters, and training a formula retriever to retrieve formulas which are then incorporated into the LLM prompts.
4.1 Dataset Split
We divided FormulaQA into training, id (in-distribution) test, and ood (out-of-distribution) test, comprising 4,608, 421 and 391 questions, respectively. We required that all formulas in the id test must appear in the training set, whereas in the ood test, each question involves at least one formula that has not been seen in the training set. This division is designed to evaluate the generalization capabilities of fine-tuned models on formulas they have not previously encountered.
4.2 LLMs
Following Kojima et al., 2022, we incorporated the phrase “Let’s think step by step” into the zero-shot prompt to guide LLMs in generating the reasoning steps. For the few-shot setting, we randomly sampled five questions from the train set to serve as examples for in-context learning. Each example includes the question text and the reasoning steps (i.e., the explanation). Examples of the prompts can be found in Appendix A.2.1.
We conducted experiments on GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters. We also evaluated on Llama2-7B Touvron et al. (2023), Qwen-7B/14B Bai et al. (2023), InternLM2-7B/20B Team (2023), including the base and chat versions of these models. We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat versions.
4.3 Fine-tuned Small Models
To examine the performance of smaller models with fine-tuning, we fine-tuned small models with sizes less than 2B parameters. We conducted experiments on mT5-base, mT5-large Xue et al. (2021), and Qwen-1.8B Bai et al. (2023) models. Given that small models have significantly weaker numerical calculation capabilities compared to LLMs Lin et al. (2020); Testolin (2023); Lu et al. (2023), we added a setting that allows the small models to perform numerical reasoning without conducting calculations themselves, instead relying on an external calculator. Specifically, one setting involves the small models directly computing the results, while the other setting requires the models to generate the formulas for the reasoning steps along with the corresponding numerical values and units of the parameters, which are then calculated with the aid of an external calculator.
4.4 Formula Retriever
We trained a formula retriever targeting the formula database on the training set. Specifically, we encoded each question using the Chinese-BERT-wwm-base Devlin et al. (2019); Cui et al. (2021) model to obtain the CLS vector of the question. Each formula in the formula database was represented by a randomly initialized vector. During training, we calculated the cosine score between the question vector and the formula vector. The retriever was then trained with in-batch negatives and contrastive learning loss Gao et al. (2021). Subsequently, for each question in the id test, we retrieved the top five formulas with the highest scores and included them in the prompt to observe the change in the performance of the LLM when provided with relevant formulas. More details can be found in Appendix A.2.2.
4.5 Implementation Details
We accessed to GPT-3.5-turbo [^4] https://platform.openai.com/docs, GLM4555https://open.bigmodel.cn/, and Qwen-max [^6] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters. For other LLMs, we conducted experiments on NVIDIA V100-32G GPUs for 7B models, and on NVIDIA A100-80G GPUs for 14B/20B models. These LLMs generated using nucleus sampling with top_p=0.8. Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0. For mT5-base and mT5-large, we set a learning rate of 5e-5 and a batch size of 32, testing the model after training for 50 epochs. For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs. For the smaller models, we also utilized nucleus sampling with top_p=0.8 for generation. In the case of the smaller models, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated. Training mT5-base, mT5-large, and Qwen-1.8B models requires 6, 12, and 12 hours respectively.
4.6 Metric
We utilized the numbat tool to evaluate the predictions generated by the model against the gold-standard answers. A prediction is deemed correct if the relative error (prediction - gold) / gold is less than 1%. We employed accuracy, which is the proportion of questions answered correctly, as our metric.
5 Experiments Results
5.1 Results of LLMs
Model Size id test ood test Avg. zero-shot CoTs GPT-3.5-turbo unknown 26.13 25.58 25.87 GLM4 >100B 65.32 65.22 65.27 Qwen-max >100B 58.67 57.80 58.25 InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs GPT-3.5-turbo unknown 32.07 29.92 31.03 GLM4 >100B 62.47 65.98 64.16 Qwen-max >100B 58.91 63.94 61.33 InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87
Table 4: Results of LLMs on FormulaQA.
The evaluation results on LLMs are shown in Table 4. GLM4 exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked Qwen-max by an average of 7.02 points in both id test and ood test. GPT-3.5-turbo, in comparison to GLM4 and Qwen-max, demonstrated inferior performance, which could be attributed to the likelihood that GLM4 and Qwen-max had access to a larger corpus of Chinese data during training, thereby enhancing their capability in Chinese. Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings. The subpar performance of Llama2 might be due to its training data being primarily in English. We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.2.3.
After incorporating few-shot examples, both GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14. However, similar performance changes were not observed in GLM4, possibly due to supervised fine-tuning and alignment with human preferences, which enhanced GLM4’s understanding of instructions but may have compromised its in-context learning ability.
The evaluation results indicate that even LLMs larger than 100B parameters can only achieve a maximum score of 65.98 on FormulaQA and the performance on LLMs with size not exceeding 20B is particularly unsatisfactory, underscoring the challenging nature of FormulaQA as an unresolved dataset.
5.2 Results of Fine-tuned Small Models
Model Size id test ood test Avg. w/o calculator mT5-base 960M 1.43 5.12 3.21 mT5-large 1.7B 19.00 10.74 15.02 Qwen-1.8B 1.8B 18.29 12.28 15.40 w/ calculator mT5-base 960M 33.49 30.18 31.90 mT5-large 1.7B 33.74 32.27 33.03 Qwen-1.8B 1.8B 64.78 54.45 59.81
Table 5: Results of fine-tuned small models on FormulaQA.
Table 5 shows the evaluation results for the fine-tuned small models, with and without the calculator. Recalling Section 4.1, because small models struggle with direct numerical calculations, we provided a calculator for additional support during evaluation. Except for mT5-base, which scored poorly, mT5-large and Qwen-1.8B performed better on the id test compared to the ood test, both with and without a calculator, indicating that out-of-domain formulas indeed affected model performance. However, despite the decrease in performance on the ood test, we observed that the small model still maintains generalization ability in formula prediction. We hope that dividing our test set into id test and ood test will be helpful for evaluating the generalization ability of fine-tuned models in future works.
After using calculators, the performance of the three small models improved by scores ranging from 18.01 to 44.41. Our findings revealed that focusing on reasoning while delegating numerical computations to the calculator could enhance the performance of small models, given their limited calculating capability. This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.
(a) An error case caused by wrong formulas.
(b) An error case caused by wrong calculation.
Figure 2: Error cases.
5.3 Results of LLMs with Formula Retriever
Model id test (zero-shot) id test (few-shot) GPT-3.5-turbo 26.13 32.07 GPT-3.5-turbo w/ FR 22.09 33.02 GLM4 65.32 62.47 GLM4 w/ FR 70.31 65.80 Qwen-max 58.67 58.91 Qwen-max w/ FR 56.06 58.91
Table 6: Results of LLMs with Formula Retriever on FormulaQA. “w/ FR” in the table means that the model is evaluated with top-5 retrieved formulas. Prompts can be found in Appendix A.2.4.
The results of LLMs experiments utilizing the formula retriever are shown in Table 6. We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts. We observed a positive enhancement on GLM4, reaching the highest performance among all models, with scores of 70.31 and 65.80 on the id test with zero-shot and id test with few-shot, respectively. This indicated that incorporating a formula retriever is indeed effective, but there is still ample room to explore how to integrate the capabilities of formula retrievers with other LLMs.
5.4 Case Study and Error Analysis
We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors. We divided the error types into two main categories: formula errors and calculation errors. Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors. We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors. We provide one example for each of the two types of errors listed in Figure 2. It could be observed that FormulaQA poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
6 Conclusion
In this paper, we introduced FormulaQA, a question answering dataset for formula-based numerical reasoning, which consists of 5,420 questions collected from Chinese junior high school physics exams. Unlike previous datasets where only a fraction of the questions relied on commonsense formulas or lacked formula annotations, we annotated the reasoning steps with formulas for each question with both manual and LLM-assisted efforts. Furthermore, we constructed a formula database by merging formulas with similar meanings, serving as an external knowledge base for subsequent retrieval-based/augmented approaches.  We evaluated FormulaQA across various sizes of LLMs, fine-tuned small models, and retrieval-augmented LLMs, demonstrating its challenging nature as an unresolved task. Additionally, FormulaQA introduces numerical reasoning questions based on domain-specific formulas, a dimension scarcely addressed in prior datasets. Our findings indicate substantial room for improvement of existing models on formula-based numerical reasoning, thus motivating future research efforts.
Limitations
In order to inspire further research, we have also translated the dataset into English unitizing LLMs. However, we have not yet accurately assessed the quality of the translated dataset. At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset. Additionally, our dataset is limited to the domain of physics. Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
Ethical Considerations
We collected examination questions from publicly available data sources, which are openly accessible and drawn from examinations across various regions in China. Our dataset focuses on the elementary physics domain and its development may contribute to natural language reasoning, automated AI systems, and physics education. We do not foresee any potential risks for now. Annotators recruited for this task were undergraduate students majoring in science and engineering from a university in China and they are all skilled in elementary physics. All annotators were compensated with course hours and academic credits.
Appendix A Appendix
A.1 Dataset Construction
A.1.1 Prompts in Formula Normalization.
The process of formula normalization is delineated into three distinct stages: the generation of natural language explanations, the extraction of the associated parameters from the explanations, and the subsequent error correction phase. The initial two stages are illustrated in Figures 3 and 4. The third stage is further splited into three specific error categories, each addressed by a dedicated prompt: input errors, where the parameters mentioned in the explanation are absent from the question; calculation errors, which occur when the calculator reports an error during the computation process; and output errors, where the final computed answer is incorrect. We provide an example here focusing on prompts for correcting calculation errors, while prompts for the other two error types can be found in our code submission. The prompts designed to correcting calculation errors are depicted in Figure 5. The entire normalization procedure employs a 6-shot prompting, an instance of which is provided herein for illustrative purposes.
A.1.2 Examples of deleted questions
The questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, wrong reference answer, and so on. Here is an example of these questions in Figure 6.
A.1.3 Semantics Based Merging for Formula Database Construction
Semantics-based merging primarily employs the LLM to comprehend formulas, ascertain if two formulas are semantically equivalent, and subsequently determine whether they can be merged into a single formula. The prompt for this procedure is illustrated in Figure 7. This approach ensures that the nuanced meanings embedded within formulas are accurately captured and evaluated for potential merging, thereby enhancing the quality of formula database.
A.2 Experiments
A.2.1 Zero-shot and Few-shot Prompts
Zero-shot and few-shot prompts are shown in Figure 8.
A.2.2 Formula Retriever
Let the number of formulas in the formula database be $N$. During training, we randomly initialized a matrix $\mathbf{F}\in\mathbb{R}^{N\times d}$, where $d$ is the hidden size and the $i$-th row in $\mathbf{F}$ represented the initial representation of the $i$-th formula in formula database. We denoted a batch of questions with a batch size of $B$ as $Q=\{q_{1},q_{2},...,q_{B}\}$. The indices of the gold-standard formulas corresponding to these $B$ questions were denoted as $L=\{l_{1},l_{2},\cdots,l_{B}\}$ (i.e. the label of $q_{i}$ is $l_{i}$, where $1\leq i\leq B$).
BERT was utilized to encode each question,
$\mathbf{h}^{i}_{cls},\mathbf{h}^{i}_{1},\cdots=\mathtt{BERT}(q_{i}),1\leq i\leq B.$ (1)
Subsequently, we took the CLS vector $\mathbf{h}^{i}_{cls}$ as the representation for the $i$-th question.
We utilized in-batch negatives and contrastive learning loss,
$\mathcal{L}=-\frac{1}{B}\sum_{1\leq i\leq B}\log\frac{\exp(\cos(\mathbf{h}^{i}_{cls},\mathbf{F}_{l_{i}}))}{\sum_{1\leq j\leq B}\exp(\cos(\mathbf{h}^{i}_{cls},\mathbf{F}_{l_{j}}))}.$ (2)
Each question might correspond to multiple correct formulas, and we ensured that the same question did not appear twice in the same batch when loading the data. Based on the implementation of Chinese-BERT-wwm-base, we tested the retrieval performance on the id test set and found that Recall@5 reached 97.69%.
A.2.3 Few-shot Experiments on the LLMs of Chat Versions
Model Size id test ood test Avg. zero-shot CoTs with LLMs of chat versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87 few-shot CoTs with LLMs of chat versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03 Llama2-chat 7B 1.97 1.00 1.50
Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaQA.
In this experiment, we could compare the performance of the same version of the model under zero-shot and few-shot settings. Results are shown in Table 7. For the chat version of the LLMs, we could observe that few-shot can effectively improve model performance, with performance improvements ranging from 1.27 to 9.18 on average across id test and ood test. Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease.
A.2.4 Prompts for LLMs with Formula Retriever
We added the formulas before each question in the few-shot setting. For the examples sampled from the training set, gold-standard formulas were added before each question. For the final question from the test set in both zero-shot and few-shot prompts, we included the top 5 retrieved formulas. The prompts are shown in Figure 9.
Figure 3: Prompt of the formula normalization stage 1.
Figure 4: Prompt of the formula normalization stage 2.
Figure 5: Prompt of the formula normalization stage 3: error correction for “calculation error”.
Figure 6: An example of deleted questions.
Figure 7: Prompt for semantics-based merging.
(a) Zero-shot prompt for LLMs.
(b) Few-shot prompt for LLMs.
Figure 8: Zero-shot and few-shot prompts for LLMs.
(a) Few-shot prompt for LLMs with formula retriever.
(b) Zero-shot prompt for LLMs with formula retriever.
Figure 9: Zero-shot and few-shot prompts for LLMs with formula retriever.

### v3

@@
-FormulaQA: A Question Answering Dataset for Formula-Based Numerical Reasoning
+FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning
 Abstract
@@
 However, existing numerical reasoning datasets seldom explicitly indicate the formulas employed during the reasoning steps.
-To bridge this gap, we propose a question answering dataset for formula-based numerical reasoning called FormulaQA, from junior high school physics examinations.
-We further conduct evaluations on LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thoughts methods and we explored the approach of using retrieval-augmented LLMs when providing an external formula database.
-We also fine-tune on smaller models with size not exceeding 2B.
-Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaQA.
-FormulaQA: A Question Answering Dataset for
-Formula-Based Numerical Reasoning
-Xiao Li and Sichen Liu and Bolin Zhu and Yin Zhu and Yiwei Liu and Gong Cheng State Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China {xiaoli.nju,sichenliu,bolinzhu,yinzhu,ywliu}@smail.nju.edu.cn gcheng@nju.edu.cn
+To bridge this gap, we construct a dataset for formula-based numerical reasoning called FormulaReasoning, which consists of 5,420 reasoning-based questions.
+We employ it to conduct evaluations of LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thought methods, and we further explore using retrieval-augmented LLMs provided with an external formula database associated with our dataset.
+We also experiment with supervised methods where we divide the reasoning process into formula generation, parameter extraction, and numerical calculation, and perform data augmentation.
+Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaReasoning.
 1 Introduction
-Dataset Math23K-F MAWPS-F FormulaQA
-# questions 23,162 2,373 5,420
-# formulas (and variants) 51 (131) 18 (46) 271 (824)
-# questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) 5,420 (100%)
-Avg. # reasoning steps 1.16 1.01 2.37
-Table 1: Statistics of Math23-F, MAWPS-F and our FormulaQA.
-Figure 1: An example taken from FormulaQA.
-All the numerical values (including units) are highlighted in red font.
-In the formulas, we marked the calculated intermediate results with purple font.
-All formulas are highlighted in blue.
-Numerical reasoning constitutes one of the significant forms within natural language reasoning Frieder et al.
-(2023).
-The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs OpenAI (2023); Touvron et al.
-(2023); Li et al.
-(2023c) and specialized datasets Wang et al.
-(2017); Dua et al.
-(2019); Amini et al.
-(2019); Cobbe et al.
-(2021a).
+Numerical reasoning constitutes one of the significant forms within natural language reasoning (Frieder et al., 2023).
+The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs (OpenAI, 2023; Touvron et al., 2023; Li et al., 2023c) and specialized datasets (Wang et al., 2017; Dua et al., 2019; Amini et al., 2019; Cobbe et al., 2021a).
 Current datasets for numerical reasoning typically include simple, commonsense numerical questions that do not reflect the complexity of real-world problems.
-These datasets do not fully address the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process.
-This issue becomes particularly evident when LLMs meet hallucination Frieder et al.
-(2023); Bang et al.
-(2023).
-Consequently, one might naturally wonder that “What knowledge could I use to guide numerical reasoning process?”.
-Formulas just represent the type of knowledge that is often overlooked in research but frequently utilized in real-life applications.
-Take a question from the GSM8K Cobbe et al.
-(2021a) dataset as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber.
+These datasets have not fully addressed the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process.
+This issue becomes particularly evident when LLMs meet hallucination (Frieder et al., 2023; Bang et al., 2023).
+Consequently, one might naturally ask “What knowledge could I use to guide numerical reasoning process?”.
+Formulas exactly represent such knowledge that has been largely overlooked in research but is frequently utilized in real-life applications.
+Figure 1: An example taken from FormulaReasoning.
+Numerical values (including units) given in the question and obtained from intermediate steps are highlighted in red and purple, respectively.
+Formulas and their elements are in blue.
+Take a question from the GSM8K (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber.
 How many bolts in total does it take?”.
 This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula.
-However, in FormulaQA, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object.
-We annotated formulas for each question in FormulaQA.
-An example of FormulaQA is shown in Figure 1.
-The formula-based feature makes FormulaQA a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems.
-In fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge.
-However, existing datasets scarcely consider the incorporation of formulas into numerical reasoning.
+However, in our FormulaReasoning dataset, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object.
 Recently, Liu et al., 2023 constructed two formula-based datasets, Math23K-F and MAWPS-F.
-However, the formulas in these datasets primarily consist of commonsense formulas (such as total_amount = unit_amount $\times$ total_number), and only 33.5% and 38.4% of the questions in each dataset, respectively, require the use of formulas.
-To address this gap, we constructed a question answering dataset for numerical reasoning that requires the use of formulas called FormulaQA.
+However, the formulas in these datasets primarily consist of commonsense formulas (such as total_amount = unit_amount $\times$ total_number), and only 33.5% and 38.4% of the questions in these datasets, respectively, require the use of formulas.
+To address this gap, we constructed a dataset for numerical reasoning that requires the use of formulas called FormulaReasoning.
+We annotated formulas for each question in FormulaReasoning.
+An example of FormulaReasoning is shown in Figure 1.111Please note that FormulaReasoning is in Chinese.
+For the convenience of understanding, we translated Chinese into English in all the examples presented in this paper.
+The formula-based feature makes FormulaReasoning a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems.
+Indeed, in fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge.
+However, existing datasets scarcely consider explicit incorporation of formulas into numerical reasoning.
+Dataset Math23K-F MAWPS-F GSM8K FormulaReasoning # questions 23,162 2,373 8,792 5,420 # formulas (and variants) 51 (131) 18 (46) 0 (0) 272 (824) # questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) N/A 5,420 (100%) Avg.
+# reasoning steps 1.16 1.01 3.59 2.37
+Table 1: Statistics of Math23-F, MAWPS-F, GSM8K and our FormulaReasoning.
 We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations.
-With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database.
+With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and a set of relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database.
 The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems.
-In Table 1, we compare FormulaQA with two existing formula-based datasets.
-In comparison to Math23K-F and MAWPS-F, FormulaQA contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas.
-Additionally, all questions in FormulaQA require the use of formulas.
+In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known GSM8K.
+In comparison to Math23K-F and MAWPS-F, FormulaReasoning contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas.
+Additionally, all questions in FormulaReasoning require the use of formulas.
 The higher average number of reasoning steps (2.37 vs.
-1.16/1.01) implies that FormulaQA is more challenging and better suited for evaluating existing models as a multi-step reasoning task.
-We evaluated LLMs ranging from 7B to >100B parameters, as well as fine-tuned smaller models such as mT5-base/large Xue et al.
-(2021) and Qwen-1.8B Bai et al.
-(2023) on FormulaQA.
-We also attempted to train an encoder for formula retrieval and experimented with retrieval-augmented generation models.
-Our evaluation findings show that the best existing models only achieve an accuracy of around 70%, which is not yet satisfactory, indicating that there is still significant room for exploration in formula-based numerical reasoning.
-Our contributions can be summarized as follows:
-  • We construct a formula-based numerical reasoning dataset FormulaQA, with comprehensive annotations for each question.
-  • We conduct evaluations on LLMs of various sizes, fine-tuned smaller models, and retrieval-augmented generative models, establishing a strong baseline for future research.
-Dataset and code are available on https://github.com/nju-websoft/FormulaQA under the Apache License 2.0.
+1.16/1.01) implies that FormulaReasoning is more challenging and better suited for evaluating existing models as a multi-step formula-based reasoning task.
+We used FormulaReasoning to evaluate LLMs ranging from 7B to >100B parameters, as well as fine-tuned models such as Qwen-1.8B (Bai et al., 2023) and ChatGLM-6B (Zeng et al., 2022) with a proposed Chain-of-Thought supervised fine-tuned method and a data augmentation method.
+We also trained an encoder for formula retrieval and experimented with retrieval-augmented generative models.
+Our empirical findings show that the best existing models only achieve an accuracy of around 74%, lagging behind an accuracy 92% of humans, indicating that there is still significant room for exploration in formula-based numerical reasoning.
+Our contributions are summarized as follows:
+  • We construct a formula-based numerical reasoning dataset FormulaReasoning, with fine-grained annotations for each question.
+As a formular knowledge-guided numerical reasoning dataset, it can be applied to tasks involving trustworthy and verifiable reasoning.
+  • We conduct evaluations on LLMs of various sizes, supervised fine-tuned models, and retrieval-augmented generative models.
+The experimental results establish a strong baseline for future research and also indicate that the task remains unresolved.
+The dataset is available on https://zenodo.org/doi/10.5281/zenodo.11408109 under the CC BY 4.0 License and our code is available on https://github.com/nju-websoft/FormulaReasoning under the Apache License 2.0.
 2 Related Work
@@
 The study of numerical reasoning in natural language has existed for several years.
-Numerous datasets, such as DROP Dua et al.
-(2019), GSM8K Cobbe et al.
-(2021b), TSQA Li et al.
-(2021) and MATH Hendrycks et al.
-(2021), have introduced natural language numerical reasoning.
+Numerous datasets, such as DROP (Dua et al., 2019), GSM8K (Cobbe et al., 2021b), TSQA (Li et al., 2021) and MATH (Hendrycks et al., 2021), have introduced natural language numerical reasoning.
 Another line of research focusing on numerical reasoning in natural language is math word problem (MWP).
-MWP tasks typically provide a short passage (i.e., the question) and require the generation of an arithmetic expression that can compute the answer.
-Representative datasets include MAWPS Koncel-Kedziorski et al.
-(2016), Math23K Wang et al.
-(2017), MathQA Amini et al.
-(2019), etc.
-The recently introduced datasets Liu et al.
-(2023) Math23K-F and MAWPS-F require formulas for only 33.5% and 38.4% of the questions, respectively, and the formulas within these datasets are all simple commonsense formulas (e.g., total_cost=unit_cost $\times$ total_number).
-Our FormulaQA dataset collects questions from junior high school physics examinations, with each question accompanied by domain-specific formulas.
-In addition, we have also annotated a formula database for FormulaQA that can serve as an external knowledge base, used to assess retrieval-augmented systems.
+MWP tasks typically provide a short passage (i.e., a question) and require the generation of an arithmetic expression that can compute an answer.
+Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc.
+The recently introduced datasets (Liu et al., 2023) Math23K-F and MAWPS-F require formulas for only 33.5% and 38.4% of the questions, respectively, and the formulas within these datasets are all simple commonsense formulas (e.g., total_cost = unit_cost $\times$ total_number).
+By contrast, our FormulaReasoning dataset collects questions from junior high school physics examinations, with every question accompanied by formulas.
+In addition, we also annotated a formula database for FormulaReasoning that can serve as an external knowledge base, used to assess retrieval-augmented systems.
 2.2 Numerical Reasoning Methods
-The methods for solving numerical reasoning have evolved from statistical approaches Hosseini et al.
-(2014); Kushman et al.
-(2014) to those based on rules and templates Shi et al.
-(2015); Wang et al.
-(2019) and further to methods based on deep learning models Gupta et al.
-(2019); Chen et al.
-(2022); Kim et al.
-(2022); Li et al.
-(2023a).
+The methods for solving numerical reasoning have evolved from statistical approaches (Hosseini et al., 2014; Kushman et al., 2014) to those based on rules and templates (Shi et al., 2015; Wang et al., 2019) and further to methods based on deep learning models (Gupta et al., 2019; Chen et al., 2022; Kim et al., 2022; Li et al., 2023a).
 In the past two years, with the rapid development of LLMs, LLMs have demonstrated strong capabilities in resolving numerical reasoning questions.
-Consequently, several methods aimed at enhancing the reasoning abilities of LLMs have been proposed, including the notable Chain of Thoughts (CoTs) method Wei et al.
-(2022), along with many subsequent variant approaches Kojima et al.
-(2022); Wang et al.
-(2022); Zhou et al.
-(2022); Li et al.
-(2023b).
-We have established representative existing methods as baselines for FormulaQA, including fine-tuning on smaller-scale models with fewer than 2 billion parameters, such as mT5-base/large, and Qwen-1.8B.
-Additionally, we have applied zero-shot and few-shot CoTs prompting methods to LLMs ranging from 7 billion to over 100 billion parameters.
-Given that existing numerical reasoning methods scarcely consider retrieval-based approaches, as the existing numerical reasoning datasets mostly rely on implicit commonsense knowledge, in FormulaQA, we have trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning.
+Consequently, several methods aimed at enhancing the reasoning abilities of LLMs have been proposed, including the notable Chain of Thoughts (CoTs) method (Wei et al., 2022), along with many subsequent variant approaches (Kojima et al., 2022; Wang et al., 2022; Zhou et al., 2022; Li et al., 2023b).
+We established representative existing methods as baselines for FormulaReasoning, including zero/few-shot CoTs prompting methods to LLMs ranging from 7B to over 100B parameters.
+We trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning.
+We also divided the reasoning process into formula generation, parameter extraction, and calculation, and used data augmentation to enhance fine-tuned models with fewer than 7B parameters.
 3 Dataset Construction
-We collected raw questions from Chinese junior high school physics examinations from 2015 to the present, filtering out those that do not require formulas for solving.
+We collected raw questions from Chinese junior high school physics examinations from 2015 to the present.
+We had a total of five postgraduate volunteer students, and they all hold a bachelor’s degree in science and engineering.
 We then annotated the reasoning steps and corresponding formulas for each question.
 This process involved a combination of manual annotation and the assistance of LLMs to improve the efficiency of annotation.
-Each question included an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all parameters within the formulas.
-Finally, we compiled all the formulas we obtained and merged those expressing the same meaning to create a formula database.
-We describe this process to construct FormulaQA in detail below.
+Each question is associated with an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all the parameters within the formulas.
+Finally, we compiled all the formulas we merged those expressing the same meaning to create a formula database.
+We describe this process to construct FormulaReasoning in detail below.
 3.1 Preprocessing
 We crawled 18,433 junior high school physics examination questions in China from 2015 to the present from public sources, including only those with free-text answers and excluding multiple-choice and true/false questions.
-Each original question contains a question text and an explanation text that includes the reasoning steps.
+Each raw question contains a question text and an explanation text that includes the reasoning steps.
 We eliminated questions requiring diagrams.
 Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical.
-Utilizing a rule-based approach, we extracted the final numeric answer, including its corresponding unit, from the explanation.
-It was observed that for 487 instances, the final answer had not been accurately extracted from the explanations, prompting us to perform manual computations and annotations to determine the correct final answers for these questions.
+Utilizing a regular expression-based approach, we extracted the final numeric answer, including its unit, from the explanation.
+We found that for 487 questions, the regular expressions did not return results, so we manually annotated the positions of their answers in the text explanations.
 Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
-3.2 Formula Normalization
 Original explanation.
@@
 Explanation with normalized formulas.
-Calculating the degree of temperature increase in water: [Degree of water temperature increase] = [Final temperature] - [Initial temperature] = 60 ℃ - 20 ℃ = 40 ℃.
+1.
+Calculating the temperature increase in water: [Degree of water temperature increase] = [Final temperature] - [Initial temperature] = 60 ℃ - 20 ℃ = 40 ℃.
 The degree of water temperature increase = 40 ℃.
+2.
 Calculating the heat absorbed by water: [Heat absorbed by water] = [Mass of water] $\times$ [Specific heat capacity of water] $\times$ [Degree of water temperature increase] = 50 kg $\times$ 4.2 $\times$ $10^{3}$ J/(kg·℃) $\times$ 40 ℃ = 8400000 J.
 The heat absorbed by water = 8400000 J.
+3.
 The thermal efficiency of the water heater can be obtained from: [Thermal efficiency of the water heater] = [Heat absorbed by water] / [Total electrical energy consumed] $\times$ 100% = 8400000 J / ($1\times 10^{7}$ J) * 100% = 84%.
@@
 Table 2: Original explanation and explanation with normalized formulas (highlighted in blue).
+3.2 Formula Normalization
 We found that the reasoning steps (i.e.
@@
 Some formulas mixed parameter names (e.g., “mass of water”) and symbols (e.g., “$m_{water}$”), while others simply provided calculations in numerical form without parameter names or symbols.
-In order to ensure that explanations were based on a normalized form of formulas, we then normalized the formula annotations in the explanations.
+In order to ensure that all explanations adopted a normalized form of formulas, we normalized the formula annotations in the explanations.
 An example can be found in Table 2.
 In this process, we need to identify the formulas used within the original explanations and to correct any formatting issues.
-Manually undertaking such tasks required significant effort.
-However, since the process is not open-ended, but rather structured, methodical and verifiable, we ensured the accuracy and consistency of the numerical reasoning by applying the extraction of formulas from the explanations.
-Once the formulas were identified, we could calculate each step to ensure that the results aligned with the corresponding answers.
-Any redundant formulas or missing parameters could be detected, so we could introduce the assistance of LLMs here.
-To enhance the efficiency of the annotation process, during the LLM annotation phase, we adopted a coarse-to-fine annotation approach with the help of a LLM [^1] During dataset constructions, we accessed Qwen-max via API (https://help.aliyun.com/zh/dashscope/developer-reference/quick-start).
-Qwen-max is a LLM with over 100 billion parameters and a strong capability in understanding Chinese..
+Manually undertaking such tasks would require significant effort.
+However, since the process is not open-ended, but rather structured and verifiable, we could automatically, e.g., using a LLM, extract formulas from the explanations, calculate each step, and compare the result with the given answer to ensure the accuracy of this normalization process.
+Specifically, to enhance the efficiency of the annotation, we adopted a coarse-to-fine annotation approach with the help of a LLM [^2] During dataset construction, we accessed Qwen-max via API (https://help.aliyun.com/zh/dashscope/developer-reference/quick-start).
+Qwen-max is a LLM with over 100B parameters and a strong capability in Chinese..
 We first prompted the LLM in a few-shot manner to generate accurate explanations of the reasoning process.
@@
 Initially, we introduced the question along with its original explanation and the corresponding answer to guide the LLM through few-shot prompting to revise the original explanation.
-We observed that the ability of the LLM to revise explanations towards normalized explanations remains satisfactory.
+We observed that the ability of the LLM to revise explanations towards normalized explanations remained satisfactory.
 To assess the correctness of the revised explanations, we extracted formulas from these explanations and then computed the answer using the numbat tool [^3] https://numbat.dev.
-Numbat is designed for scientific computations with support for physical dimensions and units..
+Numbat is designed for scientific computations with support for physical units..
 In addition to providing explanations, we also required the LLM to present the values, symbols, and units of each parameter in the formulas in the form of a table.
 An example is shown in Figure 1.
-At this stage, we used a program to check the correctness of the formula format in the explanations, including whether there were omissions in parameter names, parameter symbols, and corresponding units and these issues are all correctable.
+At this stage, we checked the correctness of the formula format in the explanations by automatic rules, including whether there were omissions in parameter names, parameter symbols, or corresponding units, and these issues were all correctable.
 Therefore, if our program detected that the LLM had not successfully generated an accurate normalized explanation, we used few-shot prompting to identify and correct these specific errors.
@@
 Some examples of these questions can be found in Appendix A.1.2.
-These questions have been removed from our dataset.
+These questions were removed from our dataset.
 Following this step, our dataset contains a remaining total of 5,420 questions.
 3.3 Formula Database Construction
-Our next step was to construct a unified formula repository for the entire dataset.
+Step # Formulas Before merging 12,906 After symbolic rules based merging 1,163 After semantic-based merging 439 After manual review and error correction 272
+Table 3: Changes in the number of formulas after each merging step.
+Our next step was to construct a unified formula database for the entire dataset.
 Given that parameters in the same formula can be expressed differently across various problem contexts, for instance, the two formulas “[weight of water] = [mass of water] * [gravitational acceleration]” and “[weight] = [mass] * [gravitational acceleration]” both calculate the weight of an object, we need to merge these formulas into a single representation.
-We divided the construction process of the formulas database into three steps: 1) Merge the formulas through symbolic rules.
-2) Merge the formulas through semantics based method.
+We divided the construction process of the formula database into three steps: 1) Merge the formulas through symbolic rules.
+2) Merge the formulas through semantic-based method.
 3) Manual review and error correction.
 In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
-Step # Formulas
-Before merging 12,906
-After symbolic rules based merging 1,163
-After semantics based merging 439
-After manual review and error correction 272
-Table 3: Changes in the number of formulas after each merging step.
 Symbolic rules based merging.
 In this step, we merged formulas through symbolic rules.
-Specifically, this was achieved by comparing the structure of the formulas and the symbols contained within them.
-Taking this as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}$=($b_{1}$+$c_{1}$)/$d_{1}$”, “$f_{2}:a_{2}$=($b_{2}$+$c_{2}$)/$d_{2}$” and “$f_{3}:b_{1}$=$a_{1}$*$d_{1}$-$c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by moving $b_{1}$ to the left side of the equation.
+Specifically, this was achieved by comparing the structure of the formulas and the symbols.Take the following as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}$=($b_{1}$+$c_{1}$)/$d_{1}$”, “$f_{2}:a_{2}$=($b_{2}$+$c_{2}$)/$d_{2}$” and “$f_{3}:b_{1}$=$a_{1}$*$d_{1}$-$c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by transformation.
 Moreover, in physics, certain physical quantities are conventionally represented by specific symbols.
 For example, the mass of an object is often denoted by “$m$” and the density of an object is frequently represented by the symbol “$\rho$”.
-Subscripts are then used to distinguish which specific object the physical quantity refers to, such as “$\rho_{water}$” for the density of water.
-For any two formulas, we first computed all transformations of a formula to obtain a set of all variants of the formula.
+Subscripts are then used to distinguish which specific object a physical quantity refers to, such as “$\rho_{water}$” for the density of water.
+For any two formulas, we first computed all the transformations of each formula to obtain a set of all its variants.
 Then, we compared the formula structures in the two sets to determine if two formulas were structurally equivalent.
-If they shared the same structure, we then compared whether the symbols, with subscripts removed, were identical.
+If they shared the same structure, we then compared whether their symbols, with subscripts removed, were identical.
 If they were, we considered these two formulas to be mergeable.
 When merging, we retained the parameter with the shorter length from the two.
-Semantics based merging.
+After merging based on symbolic rules, we reduced the number of formulas in the formula database from 12,906 to 1,163.
+Semantic-based merging.
 In the symbolic rules based merging process, the semantic information of the parameter names was neglected.
 This led us to perform merges grounded on the semantics of the parameter names.
-For instance, two formulas that were not merged during the symbol fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged.
+For instance, two formulas that were not merged during the symbolic fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged.
 We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, "density" and "density of water" are semantically similar).
-For formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^3] We used jieba: https://github.com/fxsjy/jieba..
+Specifically, for formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^4] We used jieba: https://github.com/fxsjy/jieba..
 When the two sets overlapped, the parameters were considered to have semantic connection, and the formulas became candidates for merging.
-Utilizing this approach, we identified a list of potential formula mergers and then consulted the LLM for a thorough evaluation of each pair.
+Utilizing this approach, we identified a set of pairs of potentially mergeable formulas and then consulted the LLM for a thorough evaluation of each pair.
 The prompts can be found in Appendix A.1.3.
+After this step, the number of formulas in the formula database was reduced to 439.
 Manual review and error correction.
-Upon completing the aforementioned merging process, we manually inspected the correctness of the merge, rectified instances where errors occurred in the merging, and manually merged formulas that were overlooked by the LLM.
+Upon completing the aforementioned merging process, we manually inspected the correctness of the results, rectified instances where errors occurred during merging, and manually merged formulas that were overlooked by the LLM.
+In this process, there were two human volunteers cross-validating the results of manual review and annotation.
+Finally, we obtained a formula database consisting of 272 formulas.
 4 Experiments Setup
-In this section, we explore several methods to address the questions within FormulaQA.
-These include prompting LLMs using zero-shot and few-shot chains of thoughts (CoTs, Wei et al., 2022; Kojima et al., 2022), fine-tuning smaller models with less than 2 billion parameters, and training a formula retriever to retrieve formulas which are then incorporated into the LLM prompts.
+In this section, we explore several methods for handling the questions within FormulaReasoning, including prompting LLMs using zero-shot and few-shot chain-of-thought (CoT, Wei et al., 2022; Kojima et al., 2022), and training a formula retriever to retrieve formulas to be incorporated into LLM prompts.
+Additionally, we employed two approaches to enhancing the reasoning abilities of fine-tuned models with fewer than 7B parameters.
+The first approach involved dividing the reasoning process into distinct steps: formula generation, parameter extraction, and numerical calculation.
+The second approach leveraged data augmentation to improve the models’ reasoning ability.
 4.1 Dataset Split
-We divided FormulaQA into training, id (in-distribution) test, and ood (out-of-distribution) test, comprising 4,608, 421 and 391 questions, respectively.
+We divided FormulaReasoning into into subsets for training, id (in-distribution) test, and ood (out-of-distribution) test, comprising 4,608, 421 and 391 questions, respectively.
 We required that all formulas in the id test must appear in the training set, whereas in the ood test, each question involves at least one formula that has not been seen in the training set.
-This division is designed to evaluate the generalization capabilities of fine-tuned models on formulas they have not previously encountered.
-4.2 LLMs
+This division is designed to evaluate the generalizability of fine-tuned models on formulas that they have not previously encountered.
+4.2 Evaluation
+4.2.1 Human Performance
+We recruited 108 students from a high school, with each student being assigned 7–8 questions.
+Each student was given 40 minutes to complete these questions.
+These questions were used as part of their in-class exercises, and at the end, each student received a gift.
+The final statistics were collected to evaluate human performance, which was consented by all the students.
+4.2.2 LLMs
 Following Kojima et al., 2022, we incorporated the phrase “Let’s think step by step” into the zero-shot prompt to guide LLMs in generating the reasoning steps.
-For the few-shot setting, we randomly sampled five questions from the train set to serve as examples for in-context learning.
+For the few-shot setting, we randomly sampled five questions from the training set to serve as examples for in-context learning.
 Each example includes the question text and the reasoning steps (i.e., the explanation).
 Examples of the prompts can be found in Appendix A.4.1.
-We conducted experiments on GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters.
-We also evaluated on Llama2-7B Touvron et al.
-(2023), Qwen-7B/14B Bai et al.
-(2023), InternLM2-7B/20B Team (2023), including the base and chat versions of these models.
-We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat versions.
-4.3 Fine-tuned Small Models
-To examine the performance of smaller models with fine-tuning, we fine-tuned small models with sizes less than 2B parameters.
-We conducted experiments on mT5-base, mT5-large Xue et al.
-(2021), and Qwen-1.8B Bai et al.
-(2023) models.
-Given that small models have significantly weaker numerical calculation capabilities compared to LLMs Lin et al.
-(2020); Testolin (2023); Lu et al.
-(2023), we added a setting that allows the small models to perform numerical reasoning without conducting calculations themselves, instead relying on an external calculator.
-Specifically, one setting involves the small models directly computing the results, while the other setting requires the models to generate the formulas for the reasoning steps along with the corresponding numerical values and units of the parameters, which are then calculated with the aid of an external calculator.
-4.4 Formula Retriever
-We trained a formula retriever targeting the formula database on the training set.
-Specifically, we encoded each question using the Chinese-BERT-wwm-base Devlin et al.
-(2019); Cui et al.
-(2021) model to obtain the CLS vector of the question.
+We conducted experiments on GPT-4-turbo, GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters.
+We also evaluated on Llama2-7B (Touvron et al., 2023), Llama3-8B (Meta, 2024), Qwen-7B/14B (Bai et al., 2023), InternLM2-7B/20B (Team, 2023), ChatGLM3-6B (Zeng et al., 2022), including the base and chat versions of these models.
+We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat or instruct versions.
+4.2.3 Formula Retriever
+We trained a formula retriever on the training set.
+Specifically, we encoded each question using the Chinese-BERT-wwm-base (Devlin et al., 2019; Cui et al., 2021) model to obtain the CLS vector of the question.
 Each formula in the formula database was represented by a randomly initialized vector.
 During training, we calculated the cosine score between the question vector and the formula vector.
-The retriever was then trained with in-batch negatives and contrastive learning loss Gao et al.
-(2021).
+The retriever was then trained with in-batch negatives and contrastive learning loss (Gao et al., 2021).
 Subsequently, for each question in the id test, we retrieved the top five formulas with the highest scores and included them in the prompt to observe the change in the performance of the LLM when provided with relevant formulas.
 More details can be found in Appendix A.4.2.
-4.5 Implementation Details
-We accessed to GPT-3.5-turbo [^4] https://platform.openai.com/docs, GLM4555https://open.bigmodel.cn/, and Qwen-max [^6] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters.
+4.2.4 Supervised Fine-tuned Models
+We found that directly prompting models possessing fewer than 7B parameters failed to produce satisfactory outcomes (for example, ChatGLM3-6B attained merely 8.99 points in a zero-shot setting).
+Therefore, we conducted supervised fine-tuning of models with fewer than 7B parameters, yet discerned that, dissimilar to larger models (such as GPT-4-turbo), smaller models did not exhibit proficient performance in numerical extraction and calculation.
+In order to augment the reasoning capabilities of smaller models, we explored two approaches for improvement.
+Chain-of-Thought Supervised Fine-Tuning (CoT-SFT)
+We decomposed the reasoning process into several steps.
+First, we instructed the model to generate the formulas required to solve the question.
+Subsequently, the parameter names within the formulas were extracted, allowing the model to retrieve the corresponding values and units from the context.
+Next, the formulas and the associated parameter values were provided to a calculator to obtain the final result.
+This approach relieved the model of the numerical calculation, allowing it to concentrate on the reasoning aspect.
+Data Augmentation (DA)
+We augmented the training dataset with the assistance of larger models.
+Firstly, we utilized a few-shot approach to prompt the LLM (Qwen-max) to generate new question-answer pairs.
+The correctness of the computation process generated by the LLM was meticulously verified using a calculator.
+Subsequently, the formulas generated by the model were extracted and normalized.
+More details could be found in Appendix A.3.1.
+4.3 Metric
+We utilized numbat to evaluate the predictions generated by the model against the gold-standard answers.
+A prediction is deemed correct if the relative error (prediction - gold) / gold is less than 1%.
+We employed accuracy, which is the proportion of questions answered correctly, as our metric.
+5 Experiments Results
+In this section, we presented the experimental results and analysis.
+Due to space constraints, the error analysis can be found in Appendix A.2 and the implementation details can be found in Appendix A.4.
+5.1 Human Performance
+In FormulaReasoning, humans achieved impressive performance, with a score of 93.49 on the id test, 90.47 on the ood test, and an average score of 92.03.
+5.2 Results of LLMs
+Model Size zero-shot CoT few-shot CoT id test ood test Avg.
+id test ood test Avg.
+GPT-4-turbo unknown 70.07 72.89 71.43 71.50 77.49 74.38 GPT-3.5-turbo unknown 26.13 25.58 25.87 32.07 29.92 31.03 GLM4 >100B 65.32 65.22 65.27 62.47 65.98 64.16 Qwen-max >100B 58.67 57.80 58.25 58.91 63.94 61.33 InternLM 20B 5.70 4.60 5.17 18.29 11.25 14.90 Qwen 14B 32.07 37.60 34.73 44.89 36.83 41.01 Llama3 8B 26.66 17.98 20.41 12.81 8.87 10.91 Llama2 7B 0.00 0.26 0.13 1.43 0.26 0.87 Qwen 7B 7.36 8.70 8.01 21.14 18.16 19.71 InternLM 7B 7.84 7.67 7.76 9.50 8.18 8.86 ChatGLM3 6B 9.36 8.62 8.99 23.89 19.95 21.92 Human - 93.49 90.47 92.03 93.49 90.47 92.03
+Table 4: Results of LLMs with zero-shot and few-shot prompting.
+indicates that the chat or instruct version of the model was used in the zero-shot setting, while the base version of the model was used in the few-shot setting.
+The evaluation results on LLMs are shown in Table 4.
+GPT-4-turbo exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked GLM4 by an average of 6.16 points in zero-shot setting and 10.22 in few-shot setting.
+Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings.
+The subpar performance of Llama2 might be due to its pre-training data being primarily in English.
+We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.4.3.
+After incorporating few-shot examples, GPT-4-turbo, GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14.
+However, similar performance changes were not observed on GLM4, possibly due to its supervised fine-tuning and alignment with human preferences which enhanced GLM4’s understanding of instructions but probably also compromised its in-context learning ability.
+Human performance surpassed the performance of few-shot GTP-4-turbo on the id and ood tests by margins of 21.99 and 13.25 points, respectively.
+Such results demonstrated that there remained a substantial gap between the current capabilities of state-of-the-art LLMs and human performance.
+This was even more pronounced when considering smaller-scale models.
+These findings underscored the challenging nature of FormulaReasoning as an unresolved dataset, and that there was significant room for improvement in LLMs as they struggled to match human levels of reasoning.
+5.3 Results of LLMs with Formula Retriever
+Model zero-shot few-shot GLM4 65.32 62.47    + formula retriever 70.31 65.80 GPT-4-turbo 70.07 71.50    + formula retriever 68.17 67.00
+Table 5: Results of LLMs with Formula Retriever on the id test.
+The results of LLMs utilizing the formula retriever are shown in Table 5.
+We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts.
+We observed a positive enhancement on GLM4, with score increments of 4.99 and 3.33 with zero-shot and few-shot, respectively, on the id test.
+However, we observed a performance decline with GPT-4-turbo.
+Specifically, we found that the top 5 retrieved formulas often included irrelevant ones, as the number of formulas required varies for each problem.
+The presence of these extraneous formulas affected the model’s performance, indicating that there is considerable room for further research in utilizing a formula database.
+5.4 Results of Supervised Fine-tuned Models
+Model Size id test ood test Avg.
+Qwen-1.8B 1.8B 55.91 44.58 50.25    + DA 56.16 45.32 50.74    + CoT-SFT 73.65 74.38 74.00 ChatGLM-6B 6B 52.95 40.64 47.02    + DA 53.44 45.32 49.53    + CoT-SFT 74.63 73.89 74.23
+Table 6: Results of supervised fine-tuned models on FormulaReasoning.
+Table 6 shows the results for the supervised fine-tuned models, with and without CoT-SFT and DA, which were detailed in Section 4.2.4.
+In most settings, both models achieved higher scores on the id test than the ood test, yet they still exhibited considerable performance on the ood test.
+This indicates that 1) the ood formulas indeed impacted model performance and 2) the models still demonstrate generalizability.
+We hope that the division of id test and ood test will be helpful for assessing the generalization ability of fine-tuned models in future works.
+It was noteworthy that with CoT-SFT, Qwen-1.8B and ChatGLM3-6B, with a mere parameter count of 1.8B and 6B, respectively, achieved performance comparable to GPT-4-turbo (though such a comparison may not be entirely fair).
+This indicated that the incorporation of CoT-SFT and the use of calculators could significantly enhance the reasoning capabilities of small models.
+Our findings revealed that focusing on reasoning with CoT while delegating numerical calculation to a calculator could enhance the performance of small models, given their limited calculating capability.
+The assistance of LLMs for data augmentation could also enhance smaller models’ reasoning capability.
+This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.
+6 Conclusion and Limitations
+We introduced FormulaReasoning, a dataset for formula-based numerical reasoning.
+We annotated the reasoning steps with formulas for each question with both manual and LLM-assisted efforts.
+Furthermore, we constructed a formula database after merging formulas with similar meanings, serving as an external knowledge base for subsequent retrieval-based/augmented approaches.
+We evaluated FormulaReasoning across various sizes of LLMs, supervised fine-tuned models, and retrieval-augmented LLMs, demonstrating its challenging nature as an unresolved task.
+Our findings indicate substantial room for improvement of existing models on formula-based numerical reasoning, thus motivating future research efforts.
+We have also translated the dataset into English unitizing LLMs.
+However, we have not yet accurately assessed the quality of the translated dataset.
+At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset.
+Additionally, our dataset is limited to the domain of physics.
+Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
+Acknowledgments and Disclosure of Funding
+This work was supported by the CIPSC-SMP-Zhipu.AI Large Model Cross-Disciplinary Fund. We are grateful to Chao Li for his suggestions and efforts in the annotation.
+Appendix A Appendix
+A.1 Dataset Construction
+A.1.1 Prompts in Formula Normalization
+The process of formula normalization is delineated into three distinct stages: the generation of natural language explanations, the extraction of the associated parameters from the explanations, and the subsequent error correction phase.
+The initial two stages are illustrated in Figures 3 and 4.
+The third stage is further splited into three specific error categories, each addressed by a dedicated prompt: input errors, where the parameters mentioned in the explanation are absent from the question; calculation errors, which occur when the calculator reports an error during the computation process; and output errors, where the final computed answer is incorrect.
+We provide an example here focusing on prompts for correcting calculation errors, while prompts for the other two error types can be found in our code submission.
+The prompts designed to correct calculation errors are depicted in Figure 5.
+The entire normalization procedure employs a 6-shot prompting, an instance of which is provided herein for illustrative purposes.
+A.1.2 Examples of Deleted Questions
+The questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, wrong reference answer, and so on.
+Here is an example of these questions in Figure 6.
+A.1.3 Semantic-based Merging for Formula Database Construction
+Semantic-based merging primarily employs the LLM to comprehend formulas, ascertain if two formulas are semantically equivalent, and subsequently determine whether they can be merged into a single formula.
+The prompt for this procedure is illustrated in Figure 7.
+This approach ensures that the nuanced meanings embedded within formulas are accurately captured and evaluated for potential merging, thereby enhancing the quality of formula database.
+A.2 Case Study and Error Analysis
+(a) An error case caused by wrong formulas.
+(b) An error case caused by wrong calculation.
+Figure 2: Error cases.
+We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors.
+We divided the error types into two main categories: formula errors and calculation errors.
+Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors.
+We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors.
+We provide one example for each of the two types of errors listed in Figure 2.
+It could be observed that FormulaReasoning poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
+A.3 Experiments
+A.3.1 Data Augmentation (DA) for FormulaReasoning
+There have been several studies utilizing large language models (LLMs) for data augmentation (Ding et al., 2024).
+The data generated in these related works (Zheng et al., 2023; Whitehouse et al., 2023) primarily focus on daily conversations or sentiment analysis and do not require rigorous numerical calculations.
+Some research on data augmentation involving numerical calculations (Shum et al., 2023) employs LLMs to generate solutions to questions to aid in training, rather than creating complete questions.
+In contrast to these approaches, our work generates complete questions that involve numerical calculations (particularly formula calculations), along with automatic improvement and selection to ensure data quality.
+In order to enhance the capabilities of models, we use LLM to generate more data for fine-tuning. We divide the process of data generation into the following several steps.
+First, we randomly generated 17,000 prompts.
+Each prompt was obtained by stacking five question-answer pairs sampled form training set.
+At the end of the prompt, LLM was required to generate the sixth question-answer pair.
+Second, we normalized the generated formulas.
+Except for the absence of manual review, the remaining steps were consistent with those in Section 3.2.
+At last, we unitized the calculator to check whether the calculation process in the data generated by the LLM is correct, and discarded the generated data with incorrect calculation processes.
+After the above steps, we finally retained more than 2500 questions.
+We found that mixing the newly generated data into the original training set did not always bring positive improvement, perhaps because the newly generated data has not undergone manual review.
+We found that randomly selecting a small portion of the newly generated data can enable the model to have performance improvement.
+We set several different mixing ratios selected from $\{5\%,10\%,15\%,20\%,2\%,30\%,35\%,40\%\}$.
+We fine-tuned the ChatGLM-6B-base using the augmented data set.
+After training for a fixed number of steps (150k and 200k), we selected the checkpoints with the smallest loss among models of different mixing ratios.
+A.4 Implementation Details
+We accessed to GPT-4-turbo, GPT-3.5-turbo [^5] https://platform.openai.com/docs, GLM4666https://open.bigmodel.cn/, and Qwen-max [^7] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters.
 For other LLMs, we conducted experiments on NVIDIA V100-32G GPUs for 7B models, and on NVIDIA A100-80G GPUs for 14B/20B models.
@@
 For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs.
-For the smaller models, we also utilized nucleus sampling with top_p=0.8 for generation.
-In the case of the smaller models, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated.
-Training mT5-base, mT5-large, and Qwen-1.8B models requires 6, 12, and 12 hours respectively.
-4.6 Metric
-We utilized the numbat tool to evaluate the predictions generated by the model against the gold-standard answers.
-A prediction is deemed correct if the relative error (prediction - gold) / gold is less than 1%.
-We employed accuracy, which is the proportion of questions answered correctly, as our metric.
-5 Experiments Results
-5.1 Results of LLMs
-Model Size id test ood test Avg.
-zero-shot CoTs GPT-3.5-turbo unknown 26.13 25.58 25.87 GLM4 >100B 65.32 65.22 65.27 Qwen-max >100B 58.67 57.80 58.25 InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs GPT-3.5-turbo unknown 32.07 29.92 31.03 GLM4 >100B 62.47 65.98 64.16 Qwen-max >100B 58.91 63.94 61.33 InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87
-Table 4: Results of LLMs on FormulaQA.
-The evaluation results on LLMs are shown in Table 4.
-GLM4 exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked Qwen-max by an average of 7.02 points in both id test and ood test.
-GPT-3.5-turbo, in comparison to GLM4 and Qwen-max, demonstrated inferior performance, which could be attributed to the likelihood that GLM4 and Qwen-max had access to a larger corpus of Chinese data during training, thereby enhancing their capability in Chinese.
-Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings.
-The subpar performance of Llama2 might be due to its training data being primarily in English.
-We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.2.3.
-After incorporating few-shot examples, both GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14.
-However, similar performance changes were not observed in GLM4, possibly due to supervised fine-tuning and alignment with human preferences, which enhanced GLM4’s understanding of instructions but may have compromised its in-context learning ability.
-The evaluation results indicate that even LLMs larger than 100B parameters can only achieve a maximum score of 65.98 on FormulaQA and the performance on LLMs with size not exceeding 20B is particularly unsatisfactory, underscoring the challenging nature of FormulaQA as an unresolved dataset.
-5.2 Results of Fine-tuned Small Models
-Model Size id test ood test Avg.
-w/o calculator mT5-base 960M 1.43 5.12 3.21 mT5-large 1.7B 19.00 10.74 15.02 Qwen-1.8B 1.8B 18.29 12.28 15.40 w/ calculator mT5-base 960M 33.49 30.18 31.90 mT5-large 1.7B 33.74 32.27 33.03 Qwen-1.8B 1.8B 64.78 54.45 59.81
-Table 5: Results of fine-tuned small models on FormulaQA.
-Table 5 shows the evaluation results for the fine-tuned small models, with and without the calculator.
-Recalling Section 4.1, because small models struggle with direct numerical calculations, we provided a calculator for additional support during evaluation.
-Except for mT5-base, which scored poorly, mT5-large and Qwen-1.8B performed better on the id test compared to the ood test, both with and without a calculator, indicating that out-of-domain formulas indeed affected model performance.
-However, despite the decrease in performance on the ood test, we observed that the small model still maintains generalization ability in formula prediction.
-We hope that dividing our test set into id test and ood test will be helpful for evaluating the generalization ability of fine-tuned models in future works.
-After using calculators, the performance of the three small models improved by scores ranging from 18.01 to 44.41.
-Our findings revealed that focusing on reasoning while delegating numerical computations to the calculator could enhance the performance of small models, given their limited calculating capability.
-This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.
-(a) An error case caused by wrong formulas.
-(b) An error case caused by wrong calculation.
-Figure 2: Error cases.
-5.3 Results of LLMs with Formula Retriever
-Model id test (zero-shot) id test (few-shot) GPT-3.5-turbo 26.13 32.07 GPT-3.5-turbo w/ FR 22.09 33.02 GLM4 65.32 62.47 GLM4 w/ FR 70.31 65.80 Qwen-max 58.67 58.91 Qwen-max w/ FR 56.06 58.91
-Table 6: Results of LLMs with Formula Retriever on FormulaQA. “w/ FR” in the table means that the model is evaluated with top-5 retrieved formulas. Prompts can be found in Appendix A.2.4.
-The results of LLMs experiments utilizing the formula retriever are shown in Table 6.
-We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts.
-We observed a positive enhancement on GLM4, reaching the highest performance among all models, with scores of 70.31 and 65.80 on the id test with zero-shot and id test with few-shot, respectively.
-This indicated that incorporating a formula retriever is indeed effective, but there is still ample room to explore how to integrate the capabilities of formula retrievers with other LLMs.
-5.4 Case Study and Error Analysis
-We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors.
-We divided the error types into two main categories: formula errors and calculation errors.
-Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors.
-We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors.
-We provide one example for each of the two types of errors listed in Figure 2.
-It could be observed that FormulaQA poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
-6 Conclusion
-In this paper, we introduced FormulaQA, a question answering dataset for formula-based numerical reasoning, which consists of 5,420 questions collected from Chinese junior high school physics exams.
-Unlike previous datasets where only a fraction of the questions relied on commonsense formulas or lacked formula annotations, we annotated the reasoning steps with formulas for each question with both manual and LLM-assisted efforts.
-Furthermore, we constructed a formula database by merging formulas with similar meanings, serving as an external knowledge base for subsequent retrieval-based/augmented approaches.
-We evaluated FormulaQA across various sizes of LLMs, fine-tuned small models, and retrieval-augmented LLMs, demonstrating its challenging nature as an unresolved task.
-Additionally, FormulaQA introduces numerical reasoning questions based on domain-specific formulas, a dimension scarcely addressed in prior datasets.
-Our findings indicate substantial room for improvement of existing models on formula-based numerical reasoning, thus motivating future research efforts.
-Limitations
-In order to inspire further research, we have also translated the dataset into English unitizing LLMs.
-However, we have not yet accurately assessed the quality of the translated dataset.
-At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset.
-Additionally, our dataset is limited to the domain of physics.
-Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
-Ethical Considerations
-We collected examination questions from publicly available data sources, which are openly accessible and drawn from examinations across various regions in China.
-Our dataset focuses on the elementary physics domain and its development may contribute to natural language reasoning, automated AI systems, and physics education.
-We do not foresee any potential risks for now.
-Annotators recruited for this task were undergraduate students majoring in science and engineering from a university in China and they are all skilled in elementary physics.
-All annotators were compensated with course hours and academic credits.
-Appendix A Appendix
-A.1 Dataset Construction
-A.1.1 Prompts in Formula Normalization.
-The process of formula normalization is delineated into three distinct stages: the generation of natural language explanations, the extraction of the associated parameters from the explanations, and the subsequent error correction phase.
-The initial two stages are illustrated in Figures 3 and 4.
-The third stage is further splited into three specific error categories, each addressed by a dedicated prompt: input errors, where the parameters mentioned in the explanation are absent from the question; calculation errors, which occur when the calculator reports an error during the computation process; and output errors, where the final computed answer is incorrect.
-We provide an example here focusing on prompts for correcting calculation errors, while prompts for the other two error types can be found in our code submission.
-The prompts designed to correcting calculation errors are depicted in Figure 5.
-The entire normalization procedure employs a 6-shot prompting, an instance of which is provided herein for illustrative purposes.
-A.1.2 Examples of deleted questions
-The questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, wrong reference answer, and so on.
-Here is an example of these questions in Figure 6.
-A.1.3 Semantics Based Merging for Formula Database Construction
-Semantics-based merging primarily employs the LLM to comprehend formulas, ascertain if two formulas are semantically equivalent, and subsequently determine whether they can be merged into a single formula.
-The prompt for this procedure is illustrated in Figure 7.
-This approach ensures that the nuanced meanings embedded within formulas are accurately captured and evaluated for potential merging, thereby enhancing the quality of formula database.
-A.2 Experiments
-A.2.1 Zero-shot and Few-shot Prompts
+For ChatGLM3-6B, we fine-tuned with LoRA Hu et al.
+(2021) with r=8, alpha=32 and learning rate of 5e-5, batch size of 1.
+The max input length and output length are both set to 512.
+We utilized nucleus sampling with top_p=0.8 for generation.
+In the case of CoT-SFT, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated.
+Training mT5-base, mT5-large, Qwen-1.8B, ChatGLM-6B models requires 6, 12, 12 and 24 hours respectively.
+A.4.1 Zero-shot and Few-shot Prompts
 Zero-shot and few-shot prompts are shown in Figure 8.
-A.2.2 Formula Retriever
+A.4.2 Formula Retriever
 Let the number of formulas in the formula database be $N$.
@@
 Based on the implementation of Chinese-BERT-wwm-base, we tested the retrieval performance on the id test set and found that Recall@5 reached 97.69%.
-A.2.3 Few-shot Experiments on the LLMs of Chat Versions
+Models were evaluated with top-5 retrieved formulas. Prompts can be found in Appendix A.4.4. We utilized zero-shot CoTs.
+A.4.3 Few-shot Experiments on the LLMs of Chat Versions
 Model Size id test ood test Avg.
-zero-shot CoTs with LLMs of chat versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87 few-shot CoTs with LLMs of chat versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03 Llama2-chat 7B 1.97 1.00 1.50
-Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaQA.
-In this experiment, we could compare the performance of the same version of the model under zero-shot and few-shot settings.
+zero-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Llama3-instruct 8B 22.66 17.98 20.41 Llama2-chat 7B 0.00 0.26 0.13 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Llama3-base 8B 12.81 8.87 10.91 Llama2-base 7B 1.43 0.26 0.87 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 few-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Llama3-instruct 8B 6.90 6.16 6.54 Llama2-chat 7B 1.97 1.00 1.50 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03
+Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaReasoning.
+In this experiment, we compared the performance of the same version of the model under zero-shot and few-shot settings.
 Results are shown in Table 7.
 For the chat version of the LLMs, we could observe that few-shot can effectively improve model performance, with performance improvements ranging from 1.27 to 9.18 on average across id test and ood test.
-Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease.
-A.2.4 Prompts for LLMs with Formula Retriever
+Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease from base to chat versions.
+A.4.4 Prompts for LLMs with Formula Retriever
 We added the formulas before each question in the few-shot setting.
@@
 Figure 6: An example of deleted questions.
-Figure 7: Prompt for semantics-based merging.
+Figure 7: Prompt for semantic-based merging.
 (a) Zero-shot prompt for LLMs.

### v4

@@
 The application of formulas is a fundamental ability of humans when addressing numerical reasoning problems.
-However, existing numerical reasoning datasets seldom explicitly indicate the formulas employed during the reasoning steps.
+However, existing numerical reasoning datasets seldom indicate explicitly the formulas employed during the reasoning steps.
 To bridge this gap, we construct a dataset for formula-based numerical reasoning called FormulaReasoning, which consists of 5,420 reasoning-based questions.
@@
 We also experiment with supervised methods where we divide the reasoning process into formula generation, parameter extraction, and numerical calculation, and perform data augmentation.
-Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaReasoning.
+Our empirical findings underscore the significant potential for improvement in existing models when applied to our challenging, formula-driven FormulaReasoning.
 1 Introduction
@@
 This issue becomes particularly evident when LLMs meet hallucination (Frieder et al., 2023; Bang et al., 2023).
-Consequently, one might naturally ask “What knowledge could I use to guide numerical reasoning process?”.
-Formulas exactly represent such knowledge that has been largely overlooked in research but is frequently utilized in real-life applications.
+Consequently, one might naturally ask “What knowledge could be used to guide numerical reasoning process?”.
+Formulas exactly represent such knowledge that has been largely overlooked in previous research but is frequently utilized in real-life applications.
 Figure 1: An example taken from FormulaReasoning.
@@
 This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula.
-However, in our FormulaReasoning dataset, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object.
+However, in our FormulaReasoning dataset, we require domain-specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object.
 Recently, Liu et al., 2023 constructed two formula-based datasets, Math23K-F and MAWPS-F.
 However, the formulas in these datasets primarily consist of commonsense formulas (such as total_amount = unit_amount $\times$ total_number), and only 33.5% and 38.4% of the questions in these datasets, respectively, require the use of formulas.
-To address this gap, we constructed a dataset for numerical reasoning that requires the use of formulas called FormulaReasoning.
+To fill this gap, we constructed a dataset for numerical reasoning that requires the use of formulas called FormulaReasoning.
 We annotated formulas for each question in FormulaReasoning.
-An example of FormulaReasoning is shown in Figure 1.111Please note that FormulaReasoning is in Chinese.
+An example of FormulaReasoning is shown in Figure 1.111Please note that FormulaReasoning is originally in Chinese.
 For the convenience of understanding, we translated Chinese into English in all the examples presented in this paper.
@@
 However, existing datasets scarcely consider explicit incorporation of formulas into numerical reasoning.
-Dataset Math23K-F MAWPS-F GSM8K FormulaReasoning # questions 23,162 2,373 8,792 5,420 # formulas (and variants) 51 (131) 18 (46) 0 (0) 272 (824) # questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) N/A 5,420 (100%) Avg.
-# reasoning steps 1.16 1.01 3.59 2.37
-Table 1: Statistics of Math23-F, MAWPS-F, GSM8K and our FormulaReasoning.
+Table 1: Statistics of Math23-F, MAWPS-F, GSM8K, MATH and our FormulaReasoning.
+Dataset Math23K-F MAWPS-F GSM8K MATH FormulaReasoning # questions 23,162 2,373 8,792 12,500 5,420 # formulas (and variants) 51 (131) 18 (46) 0 (0) 0 (0) 272 (824) # questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) N/A N/A 5,420 (100%) Avg.
+# reasoning steps 1.16 1.01 3.59 Not Provided 2.37
 We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations.
-With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and a set of relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database.
+With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and a set of relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a consolidated formula database.
 The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems.
-In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known GSM8K.
+In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known GSM8K and MATH (Hendrycks et al., 2021).
 In comparison to Math23K-F and MAWPS-F, FormulaReasoning contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas.
@@
 1.16/1.01) implies that FormulaReasoning is more challenging and better suited for evaluating existing models as a multi-step formula-based reasoning task.
-We used FormulaReasoning to evaluate LLMs ranging from 7B to >100B parameters, as well as fine-tuned models such as Qwen-1.8B (Bai et al., 2023) and ChatGLM-6B (Zeng et al., 2022) with a proposed Chain-of-Thought supervised fine-tuned method and a data augmentation method.
+We used FormulaReasoning to evaluate LLMs ranging from 7B to $>$100B parameters, as well as fine-tuned models such as Qwen-1.8B (Bai et al., 2023) and ChatGLM3-6B (Zeng et al., 2022) with a proposed Chain-of-Thought supervised fine-tuned method and a data augmentation method.
 We also trained an encoder for formula retrieval and experimented with retrieval-augmented generative models.
-Our empirical findings show that the best existing models only achieve an accuracy of around 74%, lagging behind an accuracy 92% of humans, indicating that there is still significant room for exploration in formula-based numerical reasoning.
+Our empirical findings show that the best existing models only achieve an accuracy of around 84%, lagging behind an accuracy 92% of humans, indicating that there is still significant room for exploration in formula-based numerical reasoning.
 Our contributions are summarized as follows:
@@
 The experimental results establish a strong baseline for future research and also indicate that the task remains unresolved.
-The dataset is available on https://zenodo.org/doi/10.5281/zenodo.11408109 under the CC BY 4.0 License and our code is available on https://github.com/nju-websoft/FormulaReasoning under the Apache License 2.0.
+The dataset is available on Zenodo [^2] https://zenodo.org/doi/10.5281/zenodo.11408109 under the CC BY 4.0 License and our code is available on GitHub [^3] https://github.com/nju-websoft/FormulaReasoning under the Apache License 2.0.
 2 Related Work
@@
 Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc.
+Several works focus on specialized domains where some of the questions in their datasets require numerical reasoning.
+Examples include GeoSQA (Huang et al., 2019), which focuses on the geography domain, the STEM (Drori et al., 2023) dataset and the ScienceQA (Lu et al., 2022) which covers multiple disciplines in science and technology.
+The distinguishing feature of our FormulaReasoning is that the numerical reasoning questions within these datasets lack explicitly labeled formulas.
 The recently introduced datasets (Liu et al., 2023) Math23K-F and MAWPS-F require formulas for only 33.5% and 38.4% of the questions, respectively, and the formulas within these datasets are all simple commonsense formulas (e.g., total_cost = unit_cost $\times$ total_number).
@@
 We trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning.
-We also divided the reasoning process into formula generation, parameter extraction, and calculation, and used data augmentation to enhance fine-tuned models with fewer than 7B parameters.
+We also divided the reasoning process into formula generation, parameter extraction, and numerical calculation, and used data augmentation to enhance fine-tuned models with fewer than 7B parameters.
 3 Dataset Construction
@@
 Each question is associated with an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all the parameters within the formulas.
-Finally, we compiled all the formulas we merged those expressing the same meaning to create a formula database.
+Finally, we compiled all the formulas and we merged those expressing the same meaning to create a formula database.
 We describe this process to construct FormulaReasoning in detail below.
@@
 Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical.
-Utilizing a regular expression-based approach, we extracted the final numeric answer, including its unit, from the explanation.
+Utilizing a regular expression-based approach, we extracted the final numerical answer, including its unit, from the explanation.
 We found that for 487 questions, the regular expressions did not return results, so we manually annotated the positions of their answers in the text explanations.
 Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
+Table 2: Original explanation and explanation with normalized formulas (highlighted in blue).
 Original explanation.
@@
 Answer = 84%
-Table 2: Original explanation and explanation with normalized formulas (highlighted in blue).
 3.2 Formula Normalization
@@
 3.3 Formula Database Construction
+Table 3: Changes in the number of formulas after each merging step.
 Step # Formulas Before merging 12,906 After symbolic rules based merging 1,163 After semantic-based merging 439 After manual review and error correction 272
-Table 3: Changes in the number of formulas after each merging step.
 Our next step was to construct a unified formula database for the entire dataset.
@@
 In this step, we merged formulas through symbolic rules.
-Specifically, this was achieved by comparing the structure of the formulas and the symbols.Take the following as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}$=($b_{1}$+$c_{1}$)/$d_{1}$”, “$f_{2}:a_{2}$=($b_{2}$+$c_{2}$)/$d_{2}$” and “$f_{3}:b_{1}$=$a_{1}$*$d_{1}$-$c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by transformation.
+Specifically, this was achieved by comparing the structure of the formulas and the symbols.
+Take the following as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}=(b_{1}+c_{1})/d_{1}$”, “$f_{2}:a_{2}=(b_{2}+c_{2})/d_{2}$” and “$f_{3}:b_{1}=a_{1}*d_{1}-c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by transformation.
 Moreover, in physics, certain physical quantities are conventionally represented by specific symbols.
@@
 For instance, two formulas that were not merged during the symbolic fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged.
-We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, "density" and "density of water" are semantically similar).
+We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, ”density” and ”density of water” are semantically similar).
 Specifically, for formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^6] We used jieba: https://github.com/fxsjy/jieba..
@@
 This division is designed to evaluate the generalizability of fine-tuned models on formulas that they have not previously encountered.
-4.2 Evaluation
+4.2 Evaluated Methods
 4.2.1 Human Performance
@@
 Examples of the prompts can be found in Appendix A.2.2.
-We conducted experiments on GPT-4-turbo, GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters.
-We also evaluated on Llama2-7B (Touvron et al., 2023), Llama3-8B (Meta, 2024), Qwen-7B/14B (Bai et al., 2023), InternLM2-7B/20B (Team, 2023), ChatGLM3-6B (Zeng et al., 2022), including the base and chat versions of these models.
-We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat or instruct versions.
+We conducted experiments on GPT-4o, GPT-4-turbo, GPT-3.5-turbo, GLM-4-plus, GLM-4-flash (GLM et al., 2024), and Qwen-max.
+We also evaluated on Qwen2.5-7B/14B (Yang et al., 2024) and Llama3.1-8B (Meta, 2024).
 4.2.3 Formula Retriever
@@
 We found that directly prompting models possessing fewer than 7B parameters failed to produce satisfactory outcomes (for example, ChatGLM3-6B attained merely 8.99 points in a zero-shot setting).
-Therefore, we conducted supervised fine-tuning of models with fewer than 7B parameters, yet discerned that, dissimilar to larger models (such as GPT-4-turbo), smaller models did not exhibit proficient performance in numerical extraction and calculation.
+Therefore, we conducted supervised fine-tuning of models with fewer than 7B parameters, yet discerned that, dissimilar to larger models (such as GLM-4-plus), smaller models did not exhibit proficient performance in numerical extraction and calculation.
 In order to augment the reasoning capabilities of smaller models, we explored two approaches for improvement.
+We conducted experiments on Qwen-1.8B (Bai et al., 2023) and ChatGLM3-6B (Zeng et al., 2022).
 Chain-of-Thought Supervised Fine-Tuning (CoT-SFT)
@@
 Next, the formulas and the associated parameter values were provided to a calculator to obtain the final result.
-This approach relieved the model of the numerical calculation, allowing it to concentrate on the reasoning aspect.
+This approach relieved the model from numerical calculation, allowing it to concentrate on the reasoning aspect.
 Data Augmentation (DA)
 We augmented the training dataset with the assistance of larger models.
-Firstly, we utilized a few-shot approach to prompt the LLM (Qwen-max) to generate new question-answer pairs.
+Firstly, we utilized a few-shot approach to prompt a LLM (Qwen-max) to generate new question-answer pairs.
 The correctness of the computation process generated by the LLM was meticulously verified using a calculator.
@@
 We employed accuracy, which is the proportion of questions answered correctly, as our metric.
+4.4 Implementation Details
+We accessed to GPT-4o (gpt-4o-2024-08-06 version), GPT-4-turbo (gpt-4-1106-preview version), GPT-3.5-turbo (gpt-3.5-turbo-1106 version) [^7] https://platform.openai.com/docs, GLM-4-plus, GLM-4-flash [^8] https://open.bigmodel.cn/, Qwen-max and Qwen2.5-7B/14B [^9] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters.
+For Llama3.1, we conducted experiments on NVIDIA V100-32G GPUs.
+These LLMs generated using nucleus sampling with top_p=0.8.
+Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0.
+For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs.
+For ChatGLM3-6B, we fine-tuned with LoRA (Hu et al., 2021) with r=8, alpha=32 and learning rate of 5e-5, batch size of 1.
+The max input length and output length are both set to 512.
+We utilized nucleus sampling with top_p=0.8 for generation.
+In the case of CoT-SFT, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated.
+Training Qwen-1.8B, ChatGLM3-6B models required 12 and 24 hours respectively.
 5 Experiments Results
-In this section, we presented the experimental results and analysis.
-Due to space constraints, the error analysis can be found in Appendix A.2 and the implementation details can be found in Appendix A.4.
 5.1 Human Performance
@@
 5.2 Results of LLMs
+Table 4: Results of LLMs with zero-shot and few-shot prompting.
 Model Size zero-shot CoT few-shot CoT id test ood test Avg.
 id test ood test Avg.
-GPT-4-turbo unknown 70.07 72.89 71.43 71.50 77.49 74.38 GPT-3.5-turbo unknown 26.13 25.58 25.87 32.07 29.92 31.03 GLM4 >100B 65.32 65.22 65.27 62.47 65.98 64.16 Qwen-max >100B 58.67 57.80 58.25 58.91 63.94 61.33 InternLM 20B 5.70 4.60 5.17 18.29 11.25 14.90 Qwen 14B 32.07 37.60 34.73 44.89 36.83 41.01 Llama3 8B 26.66 17.98 20.41 12.81 8.87 10.91 Llama2 7B 0.00 0.26 0.13 1.43 0.26 0.87 Qwen 7B 7.36 8.70 8.01 21.14 18.16 19.71 InternLM 7B 7.84 7.67 7.76 9.50 8.18 8.86 ChatGLM3 6B 9.36 8.62 8.99 23.89 19.95 21.92 Human - 93.49 90.47 92.03 93.49 90.47 92.03
-Table 4: Results of LLMs with zero-shot and few-shot prompting.
-indicates that the chat or instruct version of the model was used in the zero-shot setting, while the base version of the model was used in the few-shot setting.
+GPT-4o unknown 77.20 72.38 74.88 76.01 73.66 74.88 GPT-4-turbo unknown 70.07 72.89 71.43 71.50 77.49 74.38 GPT-3.5-turbo unknown 26.13 25.58 25.87 32.07 29.92 31.03 GLM-4-plus $>$100B 84.32 81.07 82.76 82.90 85.68 84.24 GLM-4-flash unknown 71.50 71.87 71.68 61.76 67.01 64.29 Qwen-max $>$100B 57.24 60.10 58.62 55.82 61.38 58.50 Qwen2.5 14B 61.28 64.71 62.93 61.28 65.22 63.18 Qwen2.5 7B 42.04 43.73 42.38 59.62 65.73 62.56 Llama3.1 8B 13.06 9.74 11.46 9.74 9.72 9.73 Human - 93.49 90.47 92.03 93.49 90.47 92.03
 The evaluation results on LLMs are shown in Table 4.
-GPT-4-turbo exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked GLM4 by an average of 6.16 points in zero-shot setting and 10.22 in few-shot setting.
-Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings.
-The subpar performance of Llama2 might be due to its pre-training data being primarily in English.
-We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.4.3.
-After incorporating few-shot examples, GPT-4-turbo, GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14.
-However, similar performance changes were not observed on GLM4, possibly due to its supervised fine-tuning and alignment with human preferences which enhanced GLM4’s understanding of instructions but probably also compromised its in-context learning ability.
-Human performance surpassed the performance of few-shot GTP-4-turbo on the id and ood tests by margins of 21.99 and 13.25 points, respectively.
+GLM-4-plus exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked GPT-4o by an average of 7.88 points in zero-shot setting and 9.36 in few-shot setting.
+Among models with size not exceeding 20B, Qwen2.5-14B demonstrated commendable performance in both zero-shot and few-shot settings.
+The subpar performance of Llama3.1 might be due to its pre-training data being primarily in English.
+After incorporating few-shot examples, GPT-4-turbo, GPT-3.5-turbo, GLM-4-plus and Qwen2.5 demonstrated performance improvements, ranging from 0.25 to 20.18.
+However, similar performance changes were not observed on other LLMs.
+Surprisingly, the open-source Qwen2.5-14B model outperformed the closed-source Qwen-max model [^10] We have not yet found clear information indicating whether the closed-source Qwen-max is also based on version 2.5..
+Human performance surpassed the performance of the flagship model GLM-4-plus with zero-shot setting and few-shot setting by margins of 9.27 and 7.79 points, respectively.
 Such results demonstrated that there remained a substantial gap between the current capabilities of state-of-the-art LLMs and human performance.
@@
 These findings underscored the challenging nature of FormulaReasoning as an unresolved dataset, and that there was significant room for improvement in LLMs as they struggled to match human levels of reasoning.
+We also compared the chain of thought (CoT) and program of thought (PoT, Chen et al., 2023) methods, with the results presented in Appendix A.2.4.
+The results indicated that CoT and PoT demonstrated varying performances between different models and under different settings.
 5.3 Results of LLMs with Formula Retriever
-Model zero-shot few-shot GLM4 65.32 62.47    + formula retriever 70.31 65.80 GPT-4-turbo 70.07 71.50    + formula retriever 68.17 67.00
 Table 5: Results of LLMs with Formula Retriever on the id test.
+Model zero-shot few-shot GLM-4-flash 71.50 61.76    + formula retriever 70.55 62.95 Qwen2.5-7B 42.04 59.62    + formula retriever 52.96 63.66
 The results of LLMs utilizing the formula retriever are shown in Table 5.
 We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts.
-We observed a positive enhancement on GLM4, with score increments of 4.99 and 3.33 with zero-shot and few-shot, respectively, on the id test.
-However, we observed a performance decline with GPT-4-turbo.
+We observed a positive enhancement on Qwen2.5-7B, with score increments of 10.92 and 4.04 with zero-shot and few-shot, respectively, on the id test.
+However, we found that the performance was essentially on par on the GLM-4-flash.
 Specifically, we found that the top 5 retrieved formulas often included irrelevant ones, as the number of formulas required varies for each problem.
-The presence of these extraneous formulas affected the model’s performance, indicating that there is considerable room for further research in utilizing a formula database.
+The presence of these extraneous formulas affected the model’s performance, indicating that there is considerable room for further research in retrieving from a formula database.
 5.4 Results of Supervised Fine-tuned Models
+Table 6: Results of supervised fine-tuned models on FormulaReasoning.
 Model Size id test ood test Avg.
-Qwen-1.8B 1.8B 55.91 44.58 50.25    + DA 56.16 45.32 50.74    + CoT-SFT 73.65 74.38 74.00 ChatGLM-6B 6B 52.95 40.64 47.02    + DA 53.44 45.32 49.53    + CoT-SFT 74.63 73.89 74.23
-Table 6: Results of supervised fine-tuned models on FormulaReasoning.
+Qwen-1.8B 1.8B 55.91 44.58 50.25    + DA 56.16 45.32 50.74    + CoT-SFT 73.65 74.38 74.00 ChatGLM3-6B 6B 52.95 40.64 47.02    + DA 53.44 45.32 49.53    + CoT-SFT 74.63 73.89 74.23
 Table 6 shows the results for the supervised fine-tuned models, with and without CoT-SFT and DA, which were detailed in Section 4.2.4.
 In most settings, both models achieved higher scores on the id test than the ood test, yet they still exhibited considerable performance on the ood test.
-This indicates that 1) the ood formulas indeed impacted model performance and 2) the models still demonstrate generalizability.
-We hope that the division of id test and ood test will be helpful for assessing the generalization ability of fine-tuned models in future works.
-It was noteworthy that with CoT-SFT, Qwen-1.8B and ChatGLM3-6B, with a mere parameter count of 1.8B and 6B, respectively, achieved performance comparable to GPT-4-turbo (though such a comparison may not be entirely fair).
+This indicates that 1) the ood formulas indeed challenged model performance and 2) the models still demonstrated a certain level of generalizability.
+We hope that the division of id test and ood test will be helpful for assessing the generalization ability of fine-tuned models in future work.
+It was noteworthy that with CoT-SFT, Qwen-1.8B and ChatGLM3-6B, with a mere parameter count of 1.8B and 6B, respectively, achieved performance comparable to GPT-4o (though such a comparison may not be entirely fair).
 This indicated that the incorporation of CoT-SFT and the use of calculators could significantly enhance the reasoning capabilities of small models.
@@
 The assistance of LLMs for data augmentation could also enhance smaller models’ reasoning capability.
-This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.
+This finding provides valuable insights for future deployment of numerical reasoning systems with small models.
+5.5 Case Study and Error Analysis
+(a) An error case caused by wrong formulas.
+(b) An error case caused by wrong calculation.
+Figure 2: Error cases.
+We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors.
+We divided the error types into two main categories: formula errors and calculation errors.
+Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors.
+We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors.
+We provide one example for each of the two types of errors listed in Figure 2.
+It could be observed that FormulaReasoning poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
 6 Conclusion and Limitations
@@
 Our findings indicate substantial room for improvement of existing models on formula-based numerical reasoning, thus motivating future research efforts.
-We have also translated the dataset into English unitizing LLMs.
-However, we have not yet accurately assessed the quality of the translated dataset.
-At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset.
-Additionally, our dataset is limited to the domain of physics.
-Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
-Acknowledgments and Disclosure of Funding
-This work was supported by the CIPSC-SMP-Zhipu.AI Large Model Cross-Disciplinary Fund. We are grateful to Chao Li for his suggestions and efforts in the annotation.
+In the future work, we plan to utilize the formula knowledge from FormulaReasoning to improve the numerical reasoning capabilities of LLMs.
+Possible approaches include enhancing reasoning abilities through knowledge-driven methods, preference learning methods based on formula feedback.
+One limitation of this work is that our evaluation results reported in the paper were obtained from the original Chinese version of FormulaReasoning.
+We have employed a combination of LLM-based translation and manual review to release an English version of FormulaReasoning.
+Currently, we provide a preview English version in our GitHub repository, and we will release the official English version of FormulaReasoning after completing the manual review process.
+Another limitation is that, our dataset is limited to the domain of physics.
+Although junior high school physics is not overly complex and can be understood by most people which would benefit evaluation efforts, it is still possible to explore formula-based question answering data in other domains.
 Appendix A Appendix
@@
 This approach ensures that the nuanced meanings embedded within formulas are accurately captured and evaluated for potential merging, thereby enhancing the quality of formula database.
-A.2 Case Study and Error Analysis
-(a) An error case caused by wrong formulas.
-(b) An error case caused by wrong calculation.
-Figure 2: Error cases.
-We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors.
-We divided the error types into two main categories: formula errors and calculation errors.
-Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors.
-We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors.
-We provide one example for each of the two types of errors listed in Figure 2.
-It could be observed that FormulaReasoning poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
-A.3 Experiments
-A.3.1 Data Augmentation (DA) for FormulaReasoning
+A.2 Experiments
+A.2.1 Data Augmentation (DA) for FormulaReasoning
 There have been several studies utilizing large language models (LLMs) for data augmentation (Ding et al., 2024).
@@
 We set several different mixing ratios selected from $\{5\%,10\%,15\%,20\%,2\%,30\%,35\%,40\%\}$.
-We fine-tuned the ChatGLM-6B-base using the augmented data set.
+We fine-tuned each model using the augmented data set.
 After training for a fixed number of steps (150k and 200k), we selected the checkpoints with the smallest loss among models of different mixing ratios.
-A.4 Implementation Details
-We accessed to GPT-4-turbo, GPT-3.5-turbo [^5] https://platform.openai.com/docs, GLM4666https://open.bigmodel.cn/, and Qwen-max [^7] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters.
-For other LLMs, we conducted experiments on NVIDIA V100-32G GPUs for 7B models, and on NVIDIA A100-80G GPUs for 14B/20B models.
-These LLMs generated using nucleus sampling with top_p=0.8.
-Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0.
-For mT5-base and mT5-large, we set a learning rate of 5e-5 and a batch size of 32, testing the model after training for 50 epochs.
-For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs.
-For ChatGLM3-6B, we fine-tuned with LoRA Hu et al.
-(2021) with r=8, alpha=32 and learning rate of 5e-5, batch size of 1.
-The max input length and output length are both set to 512.
-We utilized nucleus sampling with top_p=0.8 for generation.
-In the case of CoT-SFT, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated.
-Training mT5-base, mT5-large, Qwen-1.8B, ChatGLM-6B models requires 6, 12, 12 and 24 hours respectively.
-A.4.1 Zero-shot and Few-shot Prompts
+A.2.2 Zero-shot and Few-shot Prompts
 Zero-shot and few-shot prompts are shown in Figure 8.
-A.4.2 Formula Retriever
+A.2.3 Formula Retriever
 Let the number of formulas in the formula database be $N$.
@@
 Models were evaluated with top-5 retrieved formulas. Prompts can be found in Appendix A.2.5. We utilized zero-shot CoTs.
-A.4.3 Few-shot Experiments on the LLMs of Chat Versions
-Model Size id test ood test Avg.
-zero-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Llama3-instruct 8B 22.66 17.98 20.41 Llama2-chat 7B 0.00 0.26 0.13 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Llama3-base 8B 12.81 8.87 10.91 Llama2-base 7B 1.43 0.26 0.87 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 few-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Llama3-instruct 8B 6.90 6.16 6.54 Llama2-chat 7B 1.97 1.00 1.50 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03
-Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaReasoning.
-In this experiment, we compared the performance of the same version of the model under zero-shot and few-shot settings.
+A.2.4 Comparison of CoT and PoT Prompts
+Table 7: Results of LLMs with zero-shot and few-shot chain of thought (CoT) and program of thought (PoT).
+Model zero-shot few-shot id test ood test Avg.
+id test ood test Avg.
+GPT-4o (CoT) 77.20 72.38 74.88 76.01 73.66 74.88 GPT-4o (PoT) 80.76 73.91 77.46 81.47 82.61 82.02 GLM-4-plus (CoT) 84.32 81.07 82.76 82.90 85.68 84.24 GLM-4-plus (PoT) 84.08 78.51 81.40 86.70 84.91 85.84 Human 93.49 90.47 92.03 93.49 90.47 92.03
 Results are shown in Table 7.
-For the chat version of the LLMs, we could observe that few-shot can effectively improve model performance, with performance improvements ranging from 1.27 to 9.18 on average across id test and ood test.
-Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease from base to chat versions.
-A.4.4 Prompts for LLMs with Formula Retriever
+In the PoT approach, we utilized a Python interpreter to execute the code and obtain the final results.
+We found that the performance comparison between CoT and PoT varies across models.
+GPT-4o consistently demonstrated superior performance with PoT across all settings, achieving improvements of 2.58 points on average in the zero-shot setting and 7.14 points on average in the few-shot setting.
+In contrast, GLM-4-plus showed an average decline of 1.36 points in the zero-shot setting but showed an average improvement of 1.60 points in the few-shot setting.
+The finding might be related to the code capabilities of the models.
+A.2.5 Prompts for LLMs with Formula Retriever
 We added the formulas before each question in the few-shot setting.